### v1

# Zhihui Zhu

Abstract
The maximal coding rate reduction (MCR2) objective for learning structured and compact deep representations is drawing increasing attention, especially after its recent usage in the derivation of fully explainable and highly effective deep network architectures. However, it lacks a complete theoretical justification: only the properties of its global optima are known, and its global landscape has not been studied. In this work, we give a complete characterization of the properties of all its local and global optima, as well as other types of critical points. Specifically, we show that each (local or global) maximizer of the MCR2 problem corresponds to a low-dimensional, discriminative, and diverse representation, and furthermore, each critical point of the objective is either a local maximizer or a strict saddle point. Such a favorable landscape makes MCR2 a natural choice of objective for learning diverse and discriminative representations via first-order optimization methods. To validate our theoretical findings, we conduct extensive experiments on both synthetic and real data sets.
Key words: maximal coding rate reduction, representation learning, global optimality, optimization landscape, strict saddle point
1 Introduction
1.1 Background and Motivation
In the past decade, deep learning has exhibited remarkable empirical success across a wide range of engineering and scientific applications [24], such as computer vision [14, 43], natural language processing [48, 49], and health care [11], to name a few. As argued by Bengio et al. [2], Ma et al. [31], one major factor contributing to the success of deep learning is the ability of deep networks to perform powerful nonlinear feature learning by converting the data distribution to a compact and structured representation. This representation greatly facilitates various downstream tasks, including classification [10], segmentation [21], and generation [42].
Based on the theory of data compression and optimal coding [30], Chan et al. [4], Yu et al. [56] proposed a principled and unified framework for deep learning to learn a compact and structured representation. Specifically, they proposed to maximize the difference between the coding rate of all features and the sum of coding rates of features in each class, which is referred to as maximal coding rate reduction (MCR2). This problem is presented in Problem (4) and visualized in Figure 1(a). Here, the coding rate measures the “compactness" of the features, which is interpreted as the volume of a particular set spanned by the learned features: a lower coding rate implies a more compact feature set [^1] Please refer to Chan et al. [4, Section 2.1] for more details on measuring compactness of feature sets via coding rates.. Consequently, the MCR2 objective aims to maximize the volume of the set of all features while minimizing the volumes of the sets of features from each class. Motivated by the structural similarities between deep networks and unrolled optimization schemes for sparse coding [13, 34], Chan et al. [4] constructed a new deep network based on an iterative gradient descent scheme to maximize the MCR2 objective. [^2] When performing maximization, we actually mean that we use gradient ascent. However, we write gradient descent to maintain consistency with existing optimization literature. Notably, each component of this deep network has precise optimization and geometric interpretations. Moreover, it has achieved strong empirical performance on various vision and language tasks [6, 57].
(a) Visualization of feature learning via MCR2.
(b) Ideal landscape of an MCR2 problem.
Figure 1: An illustration of the properties of MCR2. (a) The high-dimensional data $\{\bm{x}_{i}\}\subseteq\mathbb{R}^{n}$ lies on a union of low-dimensional submanifolds. The objective of MCR2 is to learn a feature mapping $f_{\bm{\Theta}}(\bm{x})\in\mathbb{R}^{d}$ such that $\bm{z}_{i}=f_{\bm{\Theta}}(\bm{x}_{i})$ for all $i$ are low-dimensional, discriminative, and diverse. (b) According to Theorem 1 and Theorem 2, the regularized MCR2 problem has a benign optimization landscape: each critical point is either a local maximizer or a strict saddle point. Furthermore, each local maximizer, just like the global maximizer, corresponds to a feature representation that consists of a family of orthogonal subspaces, as illustrated in the middle.
Although the MCR2-based approach to deep learning is conceptually “white-box" and has achieved remarkable empirical performance, its theoretical foundations have been relatively under-explored. In fact, the effective feature learning mechanism and “white-box” network architecture design based on MCR2 are direct consequences of these foundations, and understanding them will pave the way to improving model interpretability and training efficiency of deep networks. Nevertheless, a comprehensive theoretical understanding of the MCR2 problem remains lacking. In this work, we take a step towards filling this gap by studying its optimization properties. Notably, analyzing these properties, including local optimality and global landscape, of the MCR2 objective is extremely challenging. To be precise, its objective function (see Problem (4)) is highly non-concave [^3] We are maximizing the MCR2 objective. Maximizing a concave function is equivalent to minimizing a convex function. and complicated, as it involves quadratic functions and the difference between log-determinant functions. To the best of our knowledge, characterizing the local optimality and global optimization landscape of the MCR2 problem remains an open question.
1.2 Our Contributions
In this work, we study the optimization foundations of the MCR2-based approach to deep learning. Towards this goal, we characterize the local and global optimality of the regularized MCR2 problem and analyze its global optimization landscape (see Problem (5)). Our contributions can be highlighted as follows.
Characterizing the local and global optimality.
For the regularized MCR2 problem, we derive the closed-form expressions for its local and global optima for the first time. Our characterization shows that each local maximizer of the regularized MCR2 problem is within-class compressible and between-class discriminative in the sense that features from the same class belong to a low-dimensional subspace, while features from different classes belong to different orthogonal subspaces. Besides these favorable properties, each global maximizer corresponds to a maximally diverse representation, which attains the highest possible dimension in the space.
Studying the global optimization landscape.
Next, we show that the regularized MCR2 function possesses a benign global optimization landscape, despite its complicated structures. More precisely, each critical point is either a local maximizer or strict saddle point of the regularized MCR2 problem; see Figure 1(b). Consequently, any gradient-based optimization, such as (stochastic) gradient descent, with random initialization can escape saddle points and at least converge to a local maximizer efficiently.
Finally, we conduct extensive numerical experiments on synthetic data sets to validate our theoretical results. Moreover, we use the regularized MCR2 objective to train deep networks on real data sets. These experimental results constitute an application of the rigorously derived MCR2 theory to more realistic and complex deep learning problems.
Our results not only establish optimization foundations for the MCR2 problem but also yield some important implications for the MCR2-based approach to deep learning. Namely, our theoretical characterizations of local and global optimality offer a compelling explanation for the empirical observations that both deep networks constructed via gradient descent applied to the MCR2 objective and over-parameterized deep networks trained by optimizing the MCR2 objective learn low-dimensional, discriminative, and diverse representations. These results align with the motivations of Chan et al. [4], Yu et al. [56] for employing the MCR2 principle for deep learning, and elucidate the outstanding performance of MCR2-based neural networks across a wide range of vision and language tasks [6, 59]. Moreover, our results underscore the potential of MCR2-based approaches to serve as a cornerstone for future advancements in deep learning, offering a principled approach to pursuing structured and compact representations in practical applications.
1.3 Related Work
Low-dimensional structures in deep representation learning.
In the literature, it has long been believed that the role of deep networks is to learn certain (nonlinear) low-dimensional and informative representations of the data [16, 31]. For example, Papyan et al. [38] showed that the features learned by cross-entropy (CE) loss exhibit a neural collapse phenomenon during the terminal phase of training, where the features from the same class are mapped to a vector while the features from different classes are maximally linearly separable. Ansuini et al. [1], Recanatesi et al. [40] demonstrated that the dimension of the intermediate features first rapidly increases and then decreases from shallow to deep layers. Masarczyk et al. [32] concluded that the deep layers of neural networks progressively compress within-class features to learn low-dimensional features. Notably, Wang et al. [53] proposed a theoretical framework to analyze hierarchical feature learning for learning low-dimensional representations. They showed that each layer of deep linear networks progressively compresses within-class features and discriminates between-class features in classification problems.
The MCR2-based approach to deep learning.
The MCR2-based approach to deep learning for seeking structured and compact representations was first proposed by Yu et al. [56]. Notably, they provided a global optimality analysis of the MCR2 problem (4) with additional rank constraints on the feature matrix of each class. Chan et al. [4] designed a new multi-layer deep network architecture, named ReduNet, based on an iterative gradient descent scheme for maximizing the MCR2 objective. To learn self-consistent representations, Dai et al. [8] extended this approach to the closed-loop transcription (CTRL) framework, which is formulated as a max-min game to optimize a modified MCR2 objective. This game was shown to have global equilibria corresponding to compact and structured representations [37]. Recently, Yu et al. [58] showed that a transformer-like architecture named CRATE, which obtains strong empirical performance [6, 57, 59], can be naturally derived through an iterative optimization scheme for maximizing the sparse rate reduction objective, which is an adaptation to sequence data of the MCR2 objective studied in this work.
Notation.
Given a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, we use $\|\bm{A}\|$ to denote its spectral norm, $\|\bm{A}\|_{F}$ to denote its Frobenius norm, and $a_{ij}$ its $(i,j)$-th element. Given a vector $\bm{a}\in\mathbb{R}^{d}$, we use $\|\bm{a}\|$ to denote its $\ell_{2}$-norm, $a_{i}$ its $i$-th element, and $\mathrm{diag}(\bm{a})$ the diagonal matrix with $\bm{a}$ on its diagonal. Given a positive integer $n$, we denote by $[n]$ the set $\{1,\ldots,n\}$. Given a set of integers $\{n_{k}\}_{k=1}^{K}$, let $n_{\max}=\max\{n_{k}:k\in[K]\}$. Let $\mathcal{O}^{m\times n}=\left\{\bm{Z}\in\mathbb{R}^{m\times n}:\bm{Z}^{T}\bm{Z}=\bm{I}_{n}\right\}$ denote the set of all $m\times n$ orthonormal matrices.
2 Problem Setup
In this section, we first review the basic concepts of MCR2 for deep representation learning in Section 2.1, and then introduce our studied problem in Section 2.2.
2.1 An Overview of MCR2
In deep representation learning, given data $\{\bm{x}_{i}\}_{i=1}^{m}\subseteq\mathbb{R}^{n}$ from multiple (say $K$) classes, the goal is to learn neural-network representations of these samples that facilitate downstream tasks. Recent empirical studies have shown that good features can be learned for tasks such as classification or autoencoding by using heuristics to promote either the contraction of samples in the same class [41] or the contrast of samples between different classes [15, 36] during the training of neural networks. Notably, Chan et al. [4], Yu et al. [56] unified and formalized these practices and demonstrated that the MCR2 objective is an effective objective to learn within-class compressible and between-class discriminative representations of the data.
The formulation of MCR2.
In this work, we mainly consider an MCR2 objective for supervised learning problems. Specifically, let $\bm{z}_{i}=f_{\bm{\Theta}}(\bm{x}_{i})$ for all $i\in[m]$ denote the features learned via the feature mapping $f_{\bm{\Theta}}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ parameterized by $\bm{\Theta}$. For each $k\in[K]$, let $\bm{\pi}^{k}\in\{0,1\}^{m}$ be a label vector denoting membership of the samples in the $k$-th class, i.e., $\pi_{i}^{k}=1$ if sample $i$ belongs to class $k$ and $\pi_{i}^{k}=0$ otherwise for all $i\in[m]$, and $m_{k}:=\sum_{i=1}^{m}\pi_{i}^{k}$ be the number of samples in the $k$-th class.
For each $k\in[K]$, let $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$ be the matrix whose columns are the features in the $k$-th class. Without loss of generality, we reorder the samples in a class-by-class manner, so that we can write the matrix of all features as
$\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}.$ (1)
On one hand, to make features between different classes discriminative or contrastive, one can maximize the lossy coding rate of all features in $\bm{Z}$, as argued in [4, 56], as follows:
$R(\bm{Z}):=\frac{1}{2}\log\det\left(\bm{I}+\frac{d}{m\epsilon^{2}}\bm{Z}\bm{Z}^{T}\right),$ (2)
where $\epsilon>0$ is a prescribed quantization error. [^4] Here, $R(\bm{Z})$ is also known as the rate-distortion function in information theory [7], which represents the average number of binary bits needed to encode the data $\bm{Z}$. On the other hand, to make features from the same class compressible or contractive, one can minimize the average lossy coding rate of features in the $k$-th class as follows:
$R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)$ $=\frac{m_{k}}{2m}\log\det\left(\bm{I}+\frac{d}{m_{k}\epsilon^{2}}\bm{Z}_{k}\bm{Z}_{k}^{T}\right).$ (3)
Consequently, a good representation tends to maximize the difference between the coding rate for the whole and that for each class as follows:
$\max_{\bm{Z}\in\mathbb{R}^{d\times m}}\ R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)\quad\mathrm{s.t.}\ \|\bm{Z}_{k}\|_{F}^{2}=m_{k},\quad\forall k\in[K].$ (4)
This is referred to as the principle of maximal coding rate reduction in [4, 56]. It is worth mentioning that this principle can be extended to self-supervised and even unsupervised learning settings, where we learn the label vectors $\{\bm{\pi}^{k}\}_{k=1}^{K}$ during training.
2.2 The Regularized MCR2 Problem
Due to the Frobenius norm constraints, it is a tremendously difficult task to analyze Problem (4) from an optimization-theoretic perspective, as all the analysis would occur on a product of spheres instead of on Euclidean space. Therefore, we consider the Lagrangian formulation of (4). This can be viewed as a tight relaxation or even an equivalent problem of (4) whose optimal solutions agree under specific settings of the regularization parameter; see Proposition 1. Specifically, the formulation we study, referred to henceforth as the regularized MCR2 problem, is as follows:
$\max_{\bm{Z}}\ F(\bm{Z}):=R(\bm{Z})-\sum_{k=1}^{K}R_{c}(\bm{Z};\bm{\pi}^{k})-\frac{\lambda}{2}\|\bm{Z}\|_{F}^{2},$ (5)
where $\lambda>0$ is the regularization parameter. Remark that our study on this problem applies meaningfully to at least two approaches to learning deep representations using the MCR2 principle.
Applications of our formulation to deep representation learning via unrolled optimization.
The first approach, as argued by Chan et al. [4], Yu et al. [57], is to construct a new deep network architecture, i.e., ReduNet [4] or CRATE [57], based on an iterative gradient descent scheme to optimize the MCR2-type objective. In this approach, each layer of the constructed network approximates a gradient descent step to optimize the MCR2-type objective given the input representation. The key takeaway is that these networks approximately implement gradient descent directly on the representations, so our analysis of the optimization properties of the MCR2-type objective translates to explanations of the corresponding properties of the learned representations and architectures of these deep networks. In particular, our argument that the optima and optimization landscape of (5) are favorable directly translates to a justification of the correctness of learned representations of the ReduNet and a validation of its architecture design. Moreover, this study enables principled improvement of deep network architectures constructed via unrolled optimization by leveraging more advanced optimization techniques better suited for problems with benign landscapes. This can improve model interpretability and efficiency.
Applications of our formulation to deep representation learning with standard neural networks.
In the second approach, one parameterizes the feature mapping $f_{\bm{\Theta}}(\cdot)$ via standard deep neural networks such as a multi-layer perceptron or ResNet [14], and treats the MCR2-type objective like other loss functions applied to outputs of a neural network, such as mean-squared error or cross-entropy loss. Studying Problem (5) from this perspective would require us to optimize over $\bm{\Theta}$ instead of over $\bm{Z}$. This new optimization problem would be extraordinarily difficult to analyze, because modern neural networks have nonlinear interactions across many layers, so the parameters $\bm{\Theta}$ would affect the final representation $\bm{Z}$ in a complex way. Fortunately, since modern neural networks are often highly over-parameterized, they can interpolate or approximate any continuous function in the feature space [29], so we may omit these constraints by assuming the unconstrained feature model, where $\bm{z}_{i}$ for all $i\in[N]$ are treated as free optimization variables [33, 55, 62, 51]. Consequently, studying the optimization properties of Problem (5) provides valuable insights into the structures of learned representations and the efficiency of training deep networks using MCR2-type objectives.
Difficulties of analyzing Problem (5).
Although Problem (5) has no constraints, one can observe that Problem (5) is highly non-concave due to the quadratic form $\bm{Z}_{k}\bm{Z}_{k}^{T}$ and the difference of log-determinant functions. Notably, this problem shares similarities with low-rank matrix factorization problems. However, it employs the log-determinant function instead of the Frobenius norm, and the computation of the objective gradient involves matrix inverses. Therefore, from an optimization point of view, it is extremely challenging to analyze Problem (5).
3 Main Results
In this section, we first characterize the local and global optimal solutions of Problem (5) in Section 3.1, and then analyze the global landscape of the objective function in Section 3.2.
3.1 Characterization of Local and Global Optimality
Although Problem (5) is highly non-concave and involves matrix inverses in its gradient computation, we can still explicitly characterize its local and global optima as follows.
Theorem 1 (Local and global optimality).
Suppose that the number of training samples in the $k$-th class is $m_{k}>0$ for each $k\in[K]$. Given a coding precision $\epsilon>0$, if the regularization parameter satisfies
$\lambda\in\left(0,\frac{d(\sqrt{m/m_{\max}}-1)}{m(\sqrt{m/m_{\max}}+1)\epsilon^{2}}\right],$ (6)
then the following statements hold:
(i) (
Characterization of local maximizers) $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]$ is a local maximizer of Problem (5) if and only if the $k$-th block admits the following decomposition
$\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}\bm{V}_{k}^{T},$ (7)
where (a) $r_{k}=\mathrm{rank}(\bm{Z}_{k})$ satisfies $r_{k}\in[0,\min\{m_{k},d\})$ and $\sum_{k=1}^{K}r_{k}\leq\min\{m,d\}$, (b) $\bm{U}_{k}\in\mathcal{O}^{d\times r_{k}}$ satisfies $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $l\neq k$, $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$, and (c) the singular value $\overline{\sigma}_{k}$ is given in (16) for each $k\in[K]$.
(ii) (
Characterization of global maximizers) $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]$ is a global maximizer of Problem (5) if and only if (a) it satisfies the above all conditions and $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$, and (b) for all $k\neq l\in[K]$ satisfying $m_{k}<m_{l}$ and $r_{l}>0$, we have $r_{k}=\min\{m_{k},d\}$.
We defer the proof to Section 4.1 and Section D.1. In this theorem, we explicitly characterize the local and global optima of Problem (5). Intuitively, this demonstrates that the features represented by each local maximizer of Problem (5) are low-dimensional and discriminative in the sense that
(i) Within-class compressible: According to (7), at each local maximizer, the features from the same class belong to the same low-dimensional linear subspace.
(ii) Between-class discriminative: It follows from (7) and $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $k\neq l$ that, at each local maximizer, the features from different classes belong to different subspaces that are orthogonal to each other.
Moreover, the features represented by each global maximizer of Problem (5) are not only low-dimensional and discriminative but also diverse in the sense that
(iii) Maximally Diverse Representation: According to $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$, at each global maximizer, the total dimension of all features is maximized to match the highest dimension that it can achieve in the feature space.
Quality of local versus global optima.
Our above discussion explains the merits of achieving both local and global optima. At each maximizer, the representations are within-class compressible and between-class discriminative (Theorem 1 (i)). Moreover, global maximizers further satisfy that the representations are all maximally diverse (Theorem 1 (ii)(a)). If all classes were balanced, i.e., $m_{1}=\cdots=m_{K}$, then Theorem 1 (ii)(b) would not apply, and these properties would be all that Theorem 1 asserts. In this case, global optima would clearly be desired over local optima. However, in the unbalanced case, the situation is more complex, because Theorem 1 (ii)(b) would apply. It says that for global optima, the classes with the smallest numbers of samples would fill to the largest dimension possible, and the very largest classes could collapse to $\bm{0}$, an undesirable situation. A dramatic example of this is when $m_{1}>\cdots>m_{K}>d$, for then any global optimum would have $\operatorname{rank}(\bm{Z}_{K})=d$ and $\bm{Z}_{1},\dots,\bm{Z}_{K-1}$ all collapse to $\bm{0}$. Overall, in the unbalanced case, global optima may not always correspond to the best representations. In particular, local optima with more equitable rank distributions (like bigger classes span more dimensions) which are still maximally diverse (i.e., ranks of each class sum to the dimension $d$) could be preferred in applications. As demonstrated in Section 5.1, these kinds of potentially useful local optima are realized in experiments, even with unbalanced classes.
Relation between Problems (4) and (5).
Based on the characterization of global optimality in Theorem 1, we show the following proposition that establishes the relationship between the constrained MCR2 problem (4) and the regularized MCR2 problem (5) in terms of their global solutions under an appropriate choice of the regularization parameter. The proof of this result can be found in Section D.2.
Proposition 1.
Suppose that the number of training samples in each class is the same, i.e., $m_{1}=\dots=m_{K}$, and the coding precision $\epsilon>0$ satisfies
$\epsilon\leq\frac{1}{6}\sqrt{\frac{d}{m}}\exp\left(-\frac{1}{2}\right)K^{-\frac{1}{K-1}}\left(1+\frac{1}{\sqrt{K}}\right)^{-\frac{2m}{K-1}}.$ (8)
The following statements hold:
(i) If $m<d$ and the regularization parameter in Problem (5) is set as
$\lambda=\frac{\alpha}{1+\alpha}-\frac{\alpha}{1+K\alpha},$ (9)
Problems (4) and (5) have the same global solution set.
(ii) If $m\geq d$, $d/K$ is an integer, and the regularization parameter in Problem (5) is set as
$\lambda=\frac{\alpha}{1+\alpha m/d}-\frac{\alpha}{1+\alpha Km/d},$ (10)
the global solution set of Problem (4) is a subset of that of Problem (5).
According to this proposition, if $\epsilon$ and $\lambda$ are appropriately chosen for Problem (5), when $m<d$, Problems (4) and (5) are equivalent in terms of their global optimal solutions; when $m\leq d$, Problem (5) is a tight Lagrangian relaxation of Problem (4) such that the global solution set of the former contains that of the latter.
3.2 Analysis of Global Optimization Landscape
While we have characterized the local and global optimal solutions in Theorem 1, it remains unknown whether these solutions can be computed efficiently using GD to solve Problem (5), as GD may get stuck at a saddle point. Fortunately, Sun et al. [44], Lee et al. [25] showed that if a function is twice continuously differentiable and satisfies strict saddle property, i.e., each critical point is either a local minimizer or a strict saddle point [^5] We say that a critical point is a strict saddle point of Problem (5) if it has a direction with strictly positive curvature; see Definition 2. This includes classical saddle points with strictly positive curvature as well as local minimizers., GD converges to its local minimizer almost surely with random initialization. We investigate the global optimization landscape of Problem (5) by characterizing all of its critical points as follows.
Theorem 2 (Benign optimization landscape).
Suppose that the number of training samples in the $k$-th class is $m_{k}>0$ for each $k\in[K]$. Given a coding precision $\epsilon>0$, if the regularization parameter satisfies (6), it holds that any critical point $\bm{Z}$ of Problem (5) that is not a local maximizer is a strict saddle point.
We defer the proof to Section 4.2 and Section D.3. Here, we make some remarks on this theorem and also on the consequences of the results derived so far.
Differences from existing results on the MCR2 problem.
Chan et al. [4], Yu et al. [56] have characterized the global optimality of Problem (4) with Frobenius norm constraints on each $\bm{Z}_{k}$ in the UFM. However, their analysis requires an additional rank constraint on each $\bm{Z}_{k}$ and only characterizes globally optimal representations. In contrast, our analysis eliminates the need for the rank constraint, and we characterize local and global optimality in Problem (5), as well as its optimization landscape. Interestingly, we demonstrate that the features represented by each local maximizer — not just global maximizers — are also compact and structured. Furthermore, we demonstrate that the regularized MCR2 objective (5) is a strict saddle function. To the best of our knowledge, Theorems 1 and 2 constitute the first analysis of local optima and optimization landscapes for MCR2 objectives. According to Daneshmand et al. [9], Lee et al. [25], Xu et al. [54], Theorems 1 and 2 imply that low-dimensional and discriminative representations can be efficiently found by (stochastic) GD on Problem (5) from a random initialization.
Comparison to existing landscape analyses in non-convex optimization.
In recent years, there has been a growing body of literature exploring optimization landscapes of non-convex problems in machine learning and deep learning. These include low-rank matrix factorization [12, 46, 5, 60], community detection [50, 52], dictionary learning [45, 39], and deep neural networks [47, 55, 61, 62, 18, 27]. The existing analyses in the literature cannot be applied to the MCR2 problem due to its special structure, which involves the log-determinant of all features minus the sum of the log-determinant of features in each class. Our work contributes to the literature on optimization landscape analyses of non-convex problems by showing that the MCR2 problem has a benign optimization landscape. Our approach may be of interest to analyses of the landscapes of other intricate loss functions in practical applications.
4 Proofs of Main Results
In this section, we sketch the proofs of our main theorems in Section 3. The complete proofs can be found in Sections B and C of the appendix. For ease of exposition, let
$\alpha:=\frac{d}{m\epsilon^{2}},\quad\alpha_{k}:=\frac{d}{m_{k}\epsilon^{2}},\quad\forall k\in[K].$ (11)
4.1 Analysis of Optimality Conditions
Our goal in this subsection is to characterize the local and global optima of Problem (5). Towards this goal, we first provide an upper bound on the objective function $F$ in Problem (5). In particular, this upper bound is tight when the blocks $\{\bm{Z}_{k}\}_{k=1}^{K}$ are orthogonal to each other. This result is a direct consequence of [4, Lemma 10].
Lemma 1.
For any $\bm{Z}=\left[\bm{Z}_{1},\ldots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}$ with $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$, we have
$F(\bm{Z})\leq\sum_{k=1}^{K}\left(\frac{1}{2}\log\det\left(\bm{I}_{n}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}_{n}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\frac{\lambda}{2}\|\bm{Z}_{k}\|_{F}^{2}\right),$ (12)
where the equality holds if and only if $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $1\leq k\neq l\leq K$.
Next, we study the following set of critical points, which are between-class discriminative (i.e., $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$):
$\mathcal{Z}:=\left\{\bm{Z}:\nabla F(\bm{Z})=0,\ \bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0},\forall k\neq l\right\}.$ (13)
Proposition 2.
Consider the setting of Theorem 1. It holds that $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\cal Z$ if and only if each $\bm{Z}_{k}$ admits the following singular value decomposition
$\bm{Z}_{k}=\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}\bm{V}_{k}^{T},\ \tilde{\bm{\Sigma}}_{k}=\mathrm{diag}\left(\sigma_{k,1},\dots,\sigma_{k,r_{k}}\right),$ (14)
where (i) $r_{k}\in\left[0,\min\{m_{k},d\}\right)$ satisfies $\sum_{k=1}^{K}r_{k}\leq d$, (ii) $\bm{U}_{k}\in\mathcal{O}^{n\times r_{k}}$ satisfies $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $1\leq k\neq l\leq K$, $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$ for all $k\in[K]$, and (iii) the singular values satisfy
$\sigma_{k,i}\in\left\{\overline{\sigma}_{k},\underline{\sigma}_{k}\right\},\ \forall i\in[r_{k}],$ (15)
where $\eta_{k}=(\alpha_{k}-\alpha)-\lambda\left(m/m_{k}+1\right)$ and
$\overline{\sigma}_{k}=\left(\frac{\eta_{k}+\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2},\quad\underline{\sigma}_{k}=\left(\frac{\eta_{k}-\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2}.$ (16)
This proposition shows that each critical point that is between-class discriminative (i.e., $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$) exhibits a specific structure: the singular values of $\bm{Z}_{k}$ can only take on two possible values, $\overline{\sigma}_{k}$ and $\underline{\sigma}_{k}$. We will leverage this structure and further show that $\bm{Z}$ is a strict saddle if there exists a $\bm{Z}_{k}$ with a singular value $\underline{\sigma}_{k}$.
4.2 Analysis of Optimization Landscape
Our goal in this subsection is to show that the function $F$ in Problem (5) has a benign optimization landscape. Towards this goal, we denote the set of critical point of $F$ by
$\mathcal{X}=\left\{\bm{Z}\in\mathbb{R}^{d\times m}:\nabla F(\bm{Z})=\bm{0}\right\}.$ (17)
According to (13), we divide the critical point set $\cal X$ into two disjoint sets $\mathcal{Z}$ and $\mathcal{Z}^{c}$, i.e., $\mathcal{X}=\mathcal{Z}\cup\mathcal{Z}^{c}$, where
$\mathcal{Z}^{c}:=\left\{\bm{Z}:\nabla F(\bm{Z})=0,\ \bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0},\ \exists k\neq l\right\}.$ (18)
Moreover, according to Proposition 2, we further divide $\cal Z$ into two disjoint sets $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Here,
$\mathcal{Z}_{1}:=\mathcal{Z}\cap\left\{\bm{Z}:\sigma_{k,i}(\bm{Z}_{k})=\overline{\sigma}_{k},\forall i\in[r_{k}],k\in[K]\right\},\ \mathcal{Z}_{2}:=\mathcal{Z}\setminus\mathcal{Z}_{1},$ (19)
where $\sigma_{k,i}(\bm{Z}_{k})$ denotes the $i$-th largest singular value of $\bm{Z}_{k}$. Our first step is to show that any point belonging to $\mathcal{Z}_{1}$ is a local maximizer, while any point belonging to $\mathcal{Z}_{2}$ is a strict saddle point.
Proposition 3.
Consider the setting of Theorem 2. Suppose that $\bm{Z}\in\cal Z$. Then, the following statements hold:
(i) If $\bm{Z}_{k}$ takes the form of (14) with $\sigma_{k,i}=\overline{\sigma}_{k}$ for all $i\in[r_{k}]$ and all $k\in[K]$, i.e., $\bm{Z}\in\mathcal{Z}_{1}$, then $\bm{Z}$ is a local maximizer.
(ii) If there exists a $k\in[K]$ and $i\in[r_{k}]$ with $r_{k}\geq 1$ such that $\sigma_{k,i}=\underline{\sigma}_{k}$, i.e., $\bm{Z}\in\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point.
Next, we proceed to the second step to show that any point belonging to $\mathcal{Z}^{c}$ is a strict saddle point. It suffices to find a direction $\bm{D}\in\mathbb{R}^{d\times m}$ such that $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0$ for each $\bm{Z}\in\mathcal{Z}^{c}$ according to Definition 2.
Proposition 4.
Consider the setting of Theorem 2. If $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point and there exists $1\leq k\neq l\leq K$ such that $\bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0}$, i.e., $\bm{Z}\in\mathcal{Z}^{c}$, then $\bm{Z}$ is a strict saddle point.
With the above preparations that characterize all the critical points, we can prove Theorem 1 and Theorem 2. We refer the reader to Appendix D for the detailed proof.
5 Experimental Results
(a) Cosine similarity
(b) The number and magnitude of singular values of each class matrix.
Figure 2: Validation of theory for the MCR2 problem. (a) We visualize the heatmap of cosine similarity among learned features by GD for solving Problem (5). The lighter pixels represent lower cosine similarities between pairwise features. (b) The blue dots are plotted based on the singular values by applying SVD to the solution returned by GD, and the red line is plotted according to the closed-form solution in (7). The number of nonzero singular values in each subspace is $24,23,27,26$, respectively.
In this section, we first conduct numerical experiments on synthetic data in Section 5.1 to validate our theoretical results, and then on real-world data sets using deep neural networks in Section 5.2 to further support our theory. All codes are implemented in Python mainly using NumPy and PyTorch. All of our experiments are executed on a computing server equipped with NVIDIA A40 GPUs. Due to space limitations, we defer some implementation details and additional experimental results to Appendix E.
5.1 Validation of Theory for Solving Problem (5)
In this subsection, we employ GD for solving Problem (5) with different parameter settings. We visualize the optimization dynamics and structures of the solutions returned by GD to verify and validate Theorems 1 and 2.
Verification of Theorem 1.
In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=30,m_{2}=70,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. Then, one can verify that $\lambda$ satisfies (6). For the solution $\bm{Z}$ returned by GD, we first plot the heatmap of the cosine similarity between pairwise columns of $\bm{Z}$ in Figure 2(a). We observe that the features from different classes are orthogonal to each other, while the features from the same class are correlated. Next, we compute the singular values of $\bm{Z}_{k}$ via singular value decomposition (SVD) and plot the singular values using blue dots for each $k\in[K]$ in Figure 2(b). According to the closed-form solution (7) in Theorem 1, we also plot the theoretical bound of singular values in red in Figure 2(b). One can observe that the number of singular values of each block is respectively $24,23,27,26$, summing up to $100$, and the red line perfectly matches the blue dots. These results all provide strong support for Theorem 1.
Verification of Theorem 2.
In this experiment, we maintain the same setting as above, except that the number of samples in each class is equal. We first fix $m=200$ and vary $d\in\{40,80,120\}$, and then fix $d=50$ and vary $d\in\{100,200,400\}$ to run GD. We plot the distances between function values of the iterates to the optimal value, which is computed according to (7) in Theorem 1, against the iteration numbers in Figure 3. We observe that GD with random initialization converges to an optimal solution at a linear rate. This indicates that the MCR2 has a benign global landscape, which supports Theorem 2.
(a) Convergence for $m=200,d\in\{40,80,120\}$
(b) Convergence for $d=50,m\in\{100,200,400\}$
Figure 3: Convergence performance of GD for solving the regularized MCR2 problem. Here, the $x$-axis is number of iterations (also denoted by $t$), and $y$-axis is the function value gap $F^{t}-F^{*}$, where $F^{t}=F(\bm{Z}^{t})$ denotes the function value at the $t$-th iterate $\bm{Z}^{t}$ generated by GD, and $F^{*}$ is the optimal value of Problem (5) computed according to (7) in Theorem 1.
5.2 Training Deep Networks Using Regularized MCR2
In this subsection, we conduct numerical experiments on the image datasets MNIST [23] and CIFAR-10 [22] to provide evidence that our theory also applies to deep networks. More specifically, we employ a multi-layer perceptron network with ReLU activation as the feature mapping $\bm{z}=f_{\bm{\Theta}}(\bm{x})$ with output dimension 32 for MNIST and 128 for CIFAR-10. Then, we train the network parameters $\bm{\Theta}$ via Adam [20] by optimizing Problem (5).
Experimental setting and results.
In the experiments, we randomly sample a balanced subset with $K$ classes and $m$ samples from MNIST or CIFAR-10, where each class has the same number of samples. We set $\lambda=0.001$ and $\epsilon=0.5$. For different subsets with corresponding values of $(m,K)$, we run experiments and report the function value $\hat{F}$ obtained by training deep networks and the optimal value $F^{*}$ computed using the closed-form solution in Theorem 1 in Table 1. To verify the discriminative nature of the features obtained by training deep networks across different classes, we measure the discrimination between features belonging to different classes by computing the cosine of the principal angle [3] between the class subspaces: $s=\max\left\{\|\bm{U}_{k}^{T}\bm{U}_{l}\|:k\neq l\in[K]\right\}\in[0,1]$, where the columns of $\bm{U}_{k}\in\mathbb{R}^{d\times r_{k}}$ are the right singular vectors corresponding to the top $r_{k}$ singular values of $\bm{Z}_{k}$ defined in (14) and $r_{k}$ is its rank [^6] We estimate the rank of a matrix by rounding its “stable rank” [17]: $r_{k}=\operatorname{round}(\|\bm{Z}_{k}\|_{F}^{2}/\|\bm{Z}_{k}\|^{2})$. for each $k\in[K]$. In particular, when $s$ is smaller, the spaces spanned by each pair $\bm{Z}_{k}$ and $\bm{Z}_{l}$ for $k\neq l$ are closer to being orthogonal to each other. Then, we record the value $s$ in Table 1 in different settings. Moreover, we visualize the pairwise cosine similarities between learned features on MNIST and CIFAR-10 when $(m,K)=(1500,6)$ and $(2500,10)$ in Figure 4.
Table 1: Function value $\hat{F}$ obtained by training deep networks, the optimal value $F^{*}$ computed by our theory on subsets of MNIST or CIFAR-10, and discrimination metric $s$ of features.
MNIST $(m,K)$ $\hat{F}$ $F^{*}$ $s$
(1000, 4) 37.38 37.38 $5.9\cdot 10^{-6}$
(1500, 6) 38.96 38.96 $3.8\cdot 10^{-6}$
(2000, 8) 38.48 38.48 0.011
(2500, 10) 37.41 37.41 0.008
CIFAR-10 $(m,K)$ $\hat{F}$ $F^{*}$ $s$
(1000, 4) 215.61 215.61 0.004
(1500, 6) 229.14 229.14 0.029
(2000, 8) 230.70 230.70 0.059
(2500, 10) 228.48 228.49 0.171
We observe from Table 1 that the function value returned by training deep networks is extremely close to the global optimal value of Problem (5) and from the value $s$ and Figure 4 that the features from different classes are nearly orthogonal to each other. These observations, together with Theorems 1 and 2, indicate that Problem (5) retains its optimization properties even when $\bm{Z}$ is parameterized by a neural network. Our theoretical analysis of Problem (5) thus illustrates a qualitative picture of training deep networks with the regularized MCR2 objective.
MNIST: $m=1500,K=6$
MNIST: $m=2500,K=10$
CIFAR: $m=1500,K=6$
CIFAR: $m=2500,K=10$
Figure 4: Heatmap of cosine similarity among features produced by deep networks trained on MNIST and CIFAR-10. The darker pixels represent higher absolute cosine similarity between features.
6 Conclusion
In this work, we provided a complete characterization of the global landscape of the MCR2 objective, a highly nonconcave and nonlinear function used for representation learning. We characterized all critical points, including the local and global optima, of the MCR2 objective, and showed that — surprisingly — it has a benign global optimization landscape. These characterizations provide rigorous justifications for why such an objective can be optimized well using simple algorithms such as gradient-based methods. In particular, we show that even local optima of the objective leads to geometrically meaningful representations. Our experimental results on synthetic and real-world datasets clearly support this new theoretical characterization. With the global landscape clearly revealed, our work paves the way for exploring better optimization strategies, hence better deep neural network architectures, for optimizing the MCR2 objective more efficiently and effectively. For future work, it is natural to extend our analysis to Problem (4) with deep network parameterizations. It is also interesting to study the sparse MCR2 objective, which has led to high-performance transformer-like architectures [57, 58].
Acknowledgements
The work of P.W. is supported in part by ARO YIP award W911NF1910027 and DoE award DE-SC0022186. The work of H.L. is supported in part by NSF China under Grant 12301403 and the Young Elite Scientists Sponsorship Program by CAST 2023QNRC001. The work of D.P. is supported by a UC Berkeley College of Engineering Fellowship. Y.Y. acknowledges support from the joint Simons Foundation-NSF DMS grant #2031899. Q.Q. acknowledges support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, and Amazon AWS AI Award, and a gift grant from KLA. Z.Z. acknowledges support from NSF grants CCF-2240708 and IIS-2312840. Y.M. acknowledges support from the joint Simons Foundation-NSF DMS grant #2031899, the ONR grant N00014-22-1-2102, and the University of Hong Kong.
Supplementary Material
The organization of the supplementary material is as follows: In Appendix A, we introduce preliminary setups and auxiliary results for studying the MCR2 problem. Then, we prove the technical results concerning the global optimality of Problem (5) in Appendix B and the optimization landscape of Problem (5) in Appendix C, respectively. In Appendix D, we prove the main theorems in Theorem 1 and Theorem 2. Finally, we provide more experimental setups and results in Appendix E.
Besides the notions introduced earlier, we shall use $\mathrm{BlkDiag}(\bm{X}_{1},\dots,\bm{X}_{K})$ to denote the block diagonal matrix whose diagonal blocks are $\bm{X}_{1},\dots,\bm{X}_{K}$.
Appendix A Preliminaries
In this section, we first introduce the first-order optimality condition and the concept of a strict saddle point for $F(\cdot)$ in Problem (5) in Section A.1, and finally present auxiliary results about matrix computations and properties of the log-determinant function in Section A.2. Recall that $\bm{Z}=\left[\bm{Z}_{1},\ldots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}$ with $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$ for each $k\in[K]$, and $\alpha,\alpha_{k}$ are defined in (11). To simplify our development, we write $R_{c}(\bm{Z},\bm{\pi}_{k})$ in (3) as
$R_{c}(\bm{Z},\bm{\pi}_{k}):=\frac{m_{k}}{m}R_{c}(\bm{Z}_{k}),\ \text{where}\ R_{c}(\bm{Z}_{k}):=\frac{1}{2}\log\det\left(\bm{I}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right).$ (20)
Therefore, we can write $F(\bm{Z})$ in Problem (5) into
$F(\bm{Z})=R(\bm{Z})-\sum_{k=1}^{K}\frac{m_{k}}{m}R_{c}(\bm{Z}_{k})-\frac{\lambda}{2}\|\bm{Z}\|_{F}^{2}.$ (21)
A.1 Optimality Conditions and Strict Saddle Points
To begin, we compute the gradient and Hessian (in bilinear form along a direction $\bm{D}\in\mathbb{R}^{d\times m}$) of $R(\cdot)$ in (2) as follows:
$\nabla R(\bm{Z})$ $=\alpha\bm{X}^{-1}\bm{Z},$ (22)
$\nabla^{2}R(\bm{Z})[\bm{D},\bm{D}]$ $=\alpha\langle\bm{X}^{-1},\bm{D}\bm{D}^{T}\rangle-\frac{\alpha^{2}}{2}\mathrm{Tr}\left(\bm{X}^{-1}(\bm{Z}\bm{D}^{T}+\bm{D}\bm{Z}^{T})\bm{X}^{-1}(\bm{Z}\bm{D}^{T}+\bm{D}\bm{Z}^{T})\right),$ (23)
where $\bm{X}:=\bm{I}_{n}+\alpha\bm{Z}\bm{Z}^{T}$ and $\alpha$ is defined in (11). Note that we can compute the gradient and Hessian of $R_{c}(\cdot)$ in (20) using the same approach. Based on the above setup, we define the first-order optimality condition of Problem (5) as follows.
Definition 1.
We say that $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point of Problem (5) if $\nabla F(\bm{Z})=\bm{0}$, i.e.,
$\alpha(\bm{I}+\alpha\bm{Z}\bm{Z}^{T})^{-1}\bm{Z}_{k}-\alpha(\bm{I}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T})^{-1}\bm{Z}_{k}-\lambda\bm{Z}_{k}=\bm{0},\ \forall k\in[K],$ (24)
where $\alpha$ and $\alpha_{k}$ are defined in (11).
According to Jin et al. [19], Lee et al. [26], we define the strict saddle point, i.e., a critical point that has a direction with strictly positive curvature [^7] Note that Problem (5) is not a minimization problem but a maximization problem., of Problem (5) as follows:
Definition 2.
Suppose that $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point of Problem (5). We say that $\bm{Z}$ is its strict saddle point if there exists a direction $\bm{D}=\left[\bm{D}_{1},\ldots,\bm{D}_{K}\right]\in\mathbb{R}^{d\times m}$ with $\bm{D}_{k}\in\mathbb{R}^{d\times m_{k}}$ such that
$\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0,$
where
$\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]=\nabla^{2}R(\bm{Z})[\bm{D},\bm{D}]-\sum_{k=1}^{K}\frac{m_{k}}{m}\nabla^{2}R_{c}(\bm{Z}_{k})[\bm{D}_{k},\bm{D}_{k}]-\lambda\|\bm{D}\|_{F}^{2}.$ (25)
Remark that for the MCR2 problem, strict saddle points include classical saddle points with strictly positive curvature as well as local minimizers.
A.2 Auxiliary Results
We provide a matrix inversion lemma, which is also known as Sherman–Morrison–Woodbury formula.
Lemma 2 (Matrix inversion lemma).
For any $\bm{Z}\in\mathbb{R}^{d\times m}$, we have
$(\bm{I}+\alpha\bm{Z}\bm{Z}^{T})^{-1}=\bm{I}-\bm{Z}\left(\frac{1}{\alpha}\bm{I}+\bm{Z}^{T}\bm{Z}\right)^{-1}\bm{Z}^{T}.$ (26)
We next present the commutative property for the log-determinant function and the upper bound for the coding rate function. We refer the reader to [4, Lemma 8 & Lemma 10] for the detailed proofs. Here, let $\bm{Z}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decompositon of $\bm{Z}\in\mathbb{R}^{d\times m}$, where $r=\mathrm{rank}(\bm{Z})\leq\min\{m,d\}$, $\bm{U}\in\mathcal{O}^{d\times r}$, $\bm{\Sigma}\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $\bm{V}\in\mathcal{O}^{m\times r}$.
Lemma 3 (Commutative property).
For any $\bm{Z}\in\mathbb{R}^{d\times m}$ and $\alpha>0$, we have
$\frac{1}{2}\log\det\left(\bm{I}+\alpha\bm{Z}\bm{Z}^{T}\right)=\frac{1}{2}\log\det\left(\bm{I}+\alpha\bm{Z}^{T}\bm{Z}\right)=\frac{1}{2}\log\det\left(\bm{I}+\alpha\bm{\Sigma}^{2}\right).$ (27)
Lemma 4.
Let $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}$. Given $\alpha>0$, it holds that
$\log\det\left(\bm{I}+\alpha\bm{Z}\bm{Z}^{T}\right)\leq\sum_{k=1}^{K}\log\det\left(\bm{I}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right),$ (28)
where the equality holds if and only if $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l\in[K]$.
Finally, we show that the objective function of Problem (5) is invariant under the block diagonal orthogonal matrices.
Lemma 5.
For any $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$, where $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for each $k\in[K]$, we have
$F(\bm{Z}\bm{O})=F(\bm{Z}),\quad\nabla F(\bm{Z}\bm{O})=\nabla F(\bm{Z})\bm{O},\quad\nabla^{2}F(\bm{Z}\bm{O})[\bm{D}\bm{O},\bm{D}\bm{O}]=\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}].$ (29)
Proof of Lemma 5.
[…]
Appendix B Proofs in Section 4.1
B.1 Proof of Lemma 1
Proof of Lemma 1.
It follows from (28) in Lemma 4 that
$\log\det\left(\bm{I}_{d}+\alpha\bm{Z}\bm{Z}^{T}\right)\leq\sum_{k=1}^{K}\log\det\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right),$ (30)
where the equality holds if and only if $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $1\leq k\neq l\leq K$. Substituting this into (5) directly yields (12). ∎
B.2 Proof of Proposition 2
Proof of Proposition 2.
Let $\bm{Z}\in\cal Z$ be arbitrary, where $\mathcal{Z}$ is defined in (13). It follows from $\bm{Z}=\left[\bm{Z}_{1},\ldots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}$ that $\sum_{k=1}^{K}r_{k}\leq d$. According to Lemma 1 and $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$ due to $\bm{Z}\in\cal Z$, we have $F(\bm{Z})=\sum_{k=1}^{K}f_{k}(\bm{Z}_{k})$, where $f_{k}:\mathbb{R}^{d\times m_{k}}\rightarrow\mathbb{R}$ takes the form of
$f_{k}(\bm{Z}_{k}):=\frac{1}{2}\log\det\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}_{d}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\frac{\lambda}{2}\|\bm{Z}_{k}\|_{F}^{2}.$ (31)
This, together with (22), yields that $\nabla F(\bm{Z})=\bm{0}$ is equivalent to
$\alpha\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)^{-1}\bm{Z}_{k}-\alpha\left(\bm{I}_{d}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)^{-1}\bm{Z}_{k}=\lambda\bm{Z}_{k},\ \forall k\in[K].$ (32)
Obviously, $\bm{Z}_{k}=\bm{0}$ is a solution of the above equation for each $k\in[K]$, which satisfies $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $l\neq k$. Now, we consider $\bm{Z}_{k}\neq\bm{0}$, and thus $1\leq r_{k}=\mathrm{rank}(\bm{Z}_{k})\leq\min\{m_{k},d\}$. Let
$\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}\bm{Q}_{k}^{T}=\begin{bmatrix}\bm{P}_{k,1}&\bm{P}_{k,2}\end{bmatrix}\begin{bmatrix}\tilde{\bm{\Sigma}}_{k}&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}\begin{bmatrix}\bm{Q}_{k,1}^{T}\\ \bm{Q}_{k,2}^{T}\end{bmatrix}$ (33)
be a singular value decomposition (SVD) of $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$, where $\tilde{\bm{\Sigma}}_{k}=\mathrm{diag}(\sigma_{k,1},\dots,\sigma_{k,r_{k}})$ with $\sigma_{k,1}\geq\dots\geq\sigma_{k,r_{k}}>0$ being positive singular values of $\bm{Z}_{k}$, $\bm{P}_{k}\in\mathcal{O}^{d}$ with $\bm{P}_{k,1}\in\mathbb{R}^{d\times r_{k}}$ and $\bm{P}_{k,2}\in\mathbb{R}^{d\times(d-r_{k})}$, and $\bm{Q}_{k}\in\mathcal{O}^{m_{k}}$ with $\bm{Q}_{k,1}\in\mathbb{R}^{m_{k}\times r_{k}}$ and $\bm{Q}_{k,2}\in\mathbb{R}^{m_{k}\times(m_{k}-r_{k})}$. Substituting this SVD into (32) yields for all $k\in[K]$,
$\alpha\bm{P}_{k}(\bm{I}_{d}+\alpha\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}\bm{Q}_{k}^{T}-\alpha\bm{P}_{k}(\bm{I}_{d}+\alpha_{k}\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}\bm{Q}_{k}^{T}=\lambda\bm{P}_{k}\bm{\Sigma}_{k}\bm{Q}_{k}^{T},$
which is equivalent to
$\alpha(\bm{I}_{d}+\alpha\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}-\alpha(\bm{I}_{d}+\alpha_{k}\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}=\lambda\bm{\Sigma}_{k}.$
Using $\bm{\Sigma}_{k}=\mathrm{BlkDiag}(\tilde{\bm{\Sigma}}_{k},\bm{0})$, we further obtain
$\alpha(\bm{I}_{r_{k}}+\alpha\tilde{\bm{\Sigma}}_{k}^{2})^{-1}\tilde{\bm{\Sigma}}_{k}-\alpha(\bm{I}_{r_{k}}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{2})^{-1}\tilde{\bm{\Sigma}}_{k}=\lambda\tilde{\bm{\Sigma}}_{k}.$
Since $\tilde{\bm{\Sigma}}_{k}$ is a diagonal matrix with diagonal entries being positive, we have for all $k\in[K]$,
$(\bm{I}_{r_{k}}+\alpha\tilde{\bm{\Sigma}}_{k}^{2})^{-1}-(\bm{I}_{r_{k}}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{2})^{-1}=\frac{\lambda}{\alpha}\bm{I}_{r_{k}}.$ (34)
This implies for each $i\in[r_{k}]$ and $k\in[K]$,
$\frac{1}{1+\alpha\sigma_{k,i}^{2}}-\frac{1}{1+\alpha_{k}\sigma_{k,i}^{2}}=\frac{\lambda}{\alpha}.$ (35)
Therefore, we obtain that $\sigma_{k,i}^{2}>0$ for each $i\in[r_{k}]$ is a positive root of the following quadratic equation with a variable $x\in\mathbb{R}$:
$\lambda\alpha_{k}x^{2}-\eta_{k}x+\frac{\lambda}{\alpha}=0,$
where
$\eta_{k}:=(\alpha_{k}-\alpha)-\lambda\left(1+\frac{\alpha_{k}}{\alpha}\right),\ \forall k\in[K].$ (36)
According to (6), one can verify that for each $k\in[K]$,
$\eta_{k}>0,\quad\eta_{k}^{2}-\frac{4\alpha_{k}}{\alpha}\lambda^{2}\geq 0.$
This yields that the above quadratic equation has positive roots as follows. For each $i\in[r_{k}]$ and $k\in[K]$, we have
$\sigma_{k,i}^{2}=\frac{\eta_{k}\pm\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}.$ (37)
Finally, using $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ and (33), we obtain $\bm{Q}_{k,1}^{T}\bm{Q}_{l,1}=\bm{0}$ for all $1\leq l\neq k\leq K$. These, together with (33), yields (14).
Conversely, suppose that each block $\bm{Z}_{k}$ of $\bm{Z}$ satisfies $\bm{Z}_{k}=\bm{0}$ or takes the form (14) for some $\bm{U}_{k}\in\mathcal{O}^{d\times r_{k}}$ satisfying $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $l\neq k$, $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$ for all $k\in[K]$, and $\sigma_{k,i}>0$ satisfying (15). We are devoted to showing $\bm{Z}\in\cal Z$. It is straightforward to verify that $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $1\leq k\neq l\leq K$. This, together with Lemma 1, implies $F(\bm{Z})=\sum_{k=1}^{K}f_{k}(\bm{Z}_{k})$. Therefore, it suffices to verify that $\nabla f_{k}(\bm{Z}_{k})=\bm{0}$ for each $k\in[K]$ in the rest of the proof. For each $k\in[K]$, if $\bm{Z}_{k}=\bm{0}$, it is obvious to verify $\nabla f(\bm{Z}_{k})=\bm{0}$. Otherwise, $\bm{Z}_{k}$ takes the form (14) for some $\bm{U}_{k}\in\mathcal{O}^{d\times r_{k}}$, $\tilde{\bm{\Sigma}}_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ satisfying (15), and $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$, where $r_{k}\geq 1$. Now, we compute for all $i\in[r_{k}]$,
$\frac{\sigma_{k,i}^{2}}{1/\alpha_{k}+\sigma_{k,i}^{2}}-\frac{\sigma_{k,i}^{2}}{1/\alpha+\sigma_{k,i}^{2}}$ $=\frac{\alpha_{k}\sigma_{k,i}^{2}}{1+\alpha_{k}\sigma_{k,i}^{2}}-\frac{\alpha\sigma_{k,i}^{2}}{1+\alpha\sigma_{k,i}^{2}}=\frac{(\alpha_{k}-\alpha)\sigma_{k,i}^{2}}{\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\left(1+\alpha\sigma_{k,i}^{2}\right)}$
$=\frac{1}{1+\alpha\sigma_{k,i}^{2}}-\frac{1}{1+\alpha_{k}\sigma_{k,i}^{2}}=\frac{\lambda}{\alpha},$ (38)
where the last equality is due to (15), (16), and (35). Then, we compute
$\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)^{-1}=\left(\bm{I}_{d}+\alpha\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}^{2}\bm{U}_{k}^{T}\right)^{-1}=\bm{I}_{d}-\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}\left(\frac{1}{\alpha}\bm{I}_{r_{k}}+\tilde{\bm{\Sigma}}_{k}^{2}\right)^{-1}\tilde{\bm{\Sigma}}_{k}\bm{U}_{k}^{T}.$ (39)
where the second equality follows from (26). This, together with (22), yields
$\nabla f_{k}(\bm{Z}_{k})$ $=\alpha\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)^{-1}\bm{Z}_{k}-\alpha\left(\bm{I}_{d}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)^{-1}\bm{Z}_{k}-\lambda\bm{Z}_{k}$
$=\alpha\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}\left(\left(\frac{1}{\alpha_{k}}\bm{I}_{r_{k}}+\tilde{\bm{\Sigma}}_{k}^{2}\right)^{-1}-\left(\frac{1}{\alpha}\bm{I}_{r_{k}}+\tilde{\bm{\Sigma}}_{k}^{2}\right)^{-1}\right)\tilde{\bm{\Sigma}}_{k}^{2}\bm{V}_{k}^{T}-\lambda\bm{Z}_{k}=\bm{0},$
where the last equality follows from (14) and (B.2). Therefore, we have $\nabla F(\bm{Z})=\bm{0}$ as desired. This, together with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$, implies $\bm{Z}\in\cal Z$. ∎
Appendix C Proofs in Section 4.2
C.1 Proof of Proposition 3
Proof of Proposition 3.
For each $\bm{Z}\in\cal Z$, it follows from Lemma 1 that
$F(\bm{Z})=\sum_{k=1}^{K}f_{k}(\bm{Z}_{k}),$ (40)
where $f_{k}$ is defined in (31). Suppose that there exists $k\in[K]$ such that $r_{k}=0$, i.e., $\bm{Z}_{k}=\bm{0}$. According to (23) and (31), we compute for any $\bm{D}_{k}\neq\bm{0}$,
$\nabla f_{k}(\bm{Z}_{k})[\bm{D}_{k},\bm{D}_{k}]=\left(\frac{\alpha}{2}-\frac{m_{k}}{2m}\alpha_{k}-\lambda\right)\|\bm{D}_{k}\|_{F}^{2}=-\lambda\|\bm{D}_{k}\|_{F}^{2}<0,$
where the second equality follows from $m_{k}\alpha_{k}/m=\alpha$ according to (11). This implies $\bm{0}$ is a local maximizer of $f_{k}(\bm{Z}_{k})$. Suppose to the contrary that $r_{k}>0$ for all $k\in[K]$. For each $\bm{Z}\in\cal Z$, using Lemma 1 with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$, (14), and (31), we have
$F(\bm{Z})$ $=\sum_{k=1}^{K}\left(\frac{1}{2}\log\det\left(\bm{I}_{n}+\alpha\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}^{2}\bm{U}_{k}^{T}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}_{n}+\alpha_{k}\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}^{2}\bm{U}_{k}^{T}\right)-\frac{\lambda}{2}\|\bm{Z}_{k}\|_{F}^{2}\right)$
$=\sum_{k=1}^{K}\left(\frac{1}{2}\log\det\left(\bm{I}+\alpha\tilde{\bm{\Sigma}}_{k}^{2}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{2}\right)-\frac{\lambda}{2}\|\tilde{\bm{\Sigma}}_{k}\|_{F}^{2}\right)$
$=\frac{1}{2}\sum_{k=1}^{K}\sum_{i=1}^{r_{k}}\left(\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)-\lambda\sigma_{k,i}^{2}\right),$ (41)
where the second equality is due to (14) and Lemma 3. For ease of exposition, let
$h_{k}(x)=\log\left(1+\alpha x\right)-\frac{m_{k}}{m}\log\left(1+\alpha_{k}x\right)-\lambda x,\ \forall k\in[k].$ (42)
Using (15), (35), (36), and (37), one can verify that $h_{k}^{\prime}(x)\leq 0$ for $x\in(0,\underline{\sigma}_{k})$, $h_{k}^{\prime}(x)\geq 0$ for $x\in[\underline{\sigma}_{k},\overline{\sigma}_{k})$, and $h_{k}^{\prime}(x)\leq 0$ for $x\in[\overline{\sigma}_{k},\infty)$ for all $k\in[K]$. This yields that $h_{k}(\underline{\sigma}_{k})$ is a local minimizer and $h(\overline{\sigma}_{k})$ is a local maximizer. This, together with (C.1) and the fact that $\bm{0}$ is a local maximizer of $f_{k}(\bm{Z}_{k})$, implies (i) and (ii).
C.2 Proof of Proposition 4
Proof of Proposition 4.
Note that $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point that satisfies (24). Suppose that $\mathrm{rank}\left(\bm{Z}\right)=r$ and $\mathrm{rank}\left(\bm{Z}_{k}\right)=r_{k}$ for all $k\in[K]$. Obviously, we have $r_{k}\leq\min\{m_{k},d\}$ for all $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq r\leq\min\{m,d\}$. Now, let $\bm{Z}\bm{Z}^{T}=\bm{Q}\bm{\Lambda}\bm{Q}^{T}$ be an eigenvalue decomposition of $\bm{Z}\bm{Z}^{T}\in\mathbb{S}_{+}^{d}$, where $\bm{Q}\in\mathcal{O}^{d\times r}$ and $\bm{\Lambda}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with diagonal entries being positive eigenvalues of $\bm{Z}\bm{Z}^{T}$. Suppose that $\bm{Z}\bm{Z}^{T}$ has $p$ distinct positive eigenvalues, where $1\leq p\leq r$. Let $\lambda_{1}>\dots>\lambda_{p}>0$ be its distinct eigenvalue values with the corresponding multiplicities being $h_{1},\dots,h_{p}\in\mathbb{N}_{+}$, respectively. Obviously, we have $\sum_{i=1}^{p}h_{i}=r$. Therefore, we write
$\bm{\Lambda}=\mathrm{BlkDiag}\left(\lambda_{1}\bm{I}_{h_{1}},\dots,\lambda_{p}\bm{I}_{h_{p}}\right),\ \bm{Q}=\begin{bmatrix}\bm{Q}_{1},\dots,\bm{Q}_{p}\end{bmatrix},$ (43)
where $\bm{Q}_{i}\in\mathcal{O}^{d\times h_{i}}$ for all $i\in[p]$.
According to Lemma 5, we can see that $\bm{Z}$ is a critical point with curvature if and only if $\bm{Z}\bm{O}$ is a critical point with the same curvature for each $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$ with $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for all $k\in[K]$. According to the SVD of $\bm{Z}_{k}$ in (33), we can take $\bm{O}_{k}=\bm{Q}_{k}$ for each $k\in[K]$. Therefore, it suffices to study $\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}$ for each $k\in[K]$. Substituting this into (24) in Definition 1 gives
$\alpha(\bm{I}+\alpha\bm{Z}\bm{Z}^{T})^{-1}\bm{P}_{k}\bm{\Sigma}_{k}-\alpha\bm{P}_{k}(\bm{I}+\alpha_{k}\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}-\lambda\bm{P}_{k}\bm{\Sigma}_{k}=\bm{0},\ \forall k\in[K].$
This is equivalent to
$\alpha(\bm{I}+\alpha\bm{Z}\bm{Z}^{T})^{-1}\bm{Z}_{k}=\bm{Z}_{k}\left(\alpha(\bm{I}+\alpha_{k}\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}+\lambda\bm{I}\right),\ \forall k\in[K].$
This yields that each column of $\bm{Z}_{k}$ is an eigenvector of $\bm{Z}$ for each $k\in[K]$. This, together with the decomposition in (43), yields that we can permute the columns of $\bm{Z}_{k}$ such that the columns belonging to the space spanned by $\bm{Q}_{i}$ are rearranged together. Let $s_{k,i}\in\mathbb{N}$ denote the number of columns of $\bm{Z}_{k}$ that belong to the space spanned by $\bm{Q}_{i}$ for each $i\in[p]$. Obviously, we have $\sum_{i=1}^{p}s_{k,i}=m_{k}$. Consequently, for each $k\in[K]$, there exists an a column permutation matrix $\bm{\Pi}_{k}\in\mathbb{R}^{m_{k}\times m_{k}}$ such that
$\bm{Z}_{k}\bm{\Pi}_{k}=\begin{bmatrix}\bm{Z}_{k}^{(1)}&\dots&\bm{Z}_{k}^{(p)}\end{bmatrix}.$ (44)
where $\bm{Q}_{i}\bm{Q}_{i}^{T}\bm{Z}_{k}^{(i)}=\bm{Z}_{k}^{(i)}\in\mathbb{R}^{d\times s_{k,i}}$. Since $\bm{Q}_{i}^{T}\bm{Q}_{j}=\bm{0}$, we have $\bm{Z}_{k}^{(i)^{T}}\bm{Z}_{k}^{(j)}=\bm{0}$ for all $i\neq j$. This, together with (20) and Lemma 4, yields
$R_{c}(\bm{Z}_{k})=\frac{m_{k}}{2m}\sum_{i=1}^{p}\log\det\left(\bm{I}_{n}+\alpha_{k}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\right).$ (45)
Moreover, let $s_{i}:=\sum_{k=1}^{K}s_{k,i}$ and
$\bm{Z}^{(i)}:=\left[\bm{Z}_{1}^{(i)}\ \dots\ \bm{Z}_{K}^{(i)}\right]\in\mathbb{R}^{d\times s_{i}},\ \forall i\in[p].$ (46)
Using this and (44), we have
$\bm{Z}\bm{Z}^{T}=\sum_{k=1}^{K}\bm{Z}_{k}\bm{Z}_{k}^{T}=\sum_{k=1}^{K}\sum_{i=1}^{p}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}=\sum_{i=1}^{p}\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}.$
This, together with (2), Lemma 4, and $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$, yields that
$R(\bm{Z})=\frac{1}{2}\sum_{i=1}^{p}\log\det\left(\bm{I}+\alpha\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}\right).$ (47)
Characterize the structure of critical points.
Now, for each $k\in[K]$ and $i\in[p]$, let $r_{k,i}=\mathrm{rank}(\bm{Z}_{k}^{(i)})$, where $r_{k,i}\leq\min\{d,s_{k,i}\}$. Moreover, let
$\bm{Z}_{k}^{(i)}=\bm{U}_{k}^{(i)}\bm{\Sigma}_{k}^{(i)}\bm{V}_{k}^{(i)^{T}}=\begin{bmatrix}\bm{U}_{k,1}^{(i)}&\bm{U}_{k,2}^{(i)}\end{bmatrix}\begin{bmatrix}\tilde{\bm{\Sigma}}_{k}^{(i)}&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}\begin{bmatrix}\bm{V}_{k,1}^{(i)^{T}}\\ \bm{V}_{k,2}^{(i)^{T}}\end{bmatrix}$ (48)
be a singular value decomposition (SVD) of $\bm{Z}_{k}^{(i)}$, where $\tilde{\bm{\Sigma}}_{k}^{(i)}\in\mathbb{R}^{r_{k,i}\times r_{k,i}}$ is a diagonal matrix with diagonal entries being positive singular values of $\bm{Z}_{k}^{(i)}$; $\bm{U}_{k}^{(i)}\in\mathcal{O}^{d}$ with $\bm{U}_{k,1}^{(i)}\in\mathbb{R}^{d\times r_{k,i}}$ and $\bm{U}_{k,2}^{(i)}\in\mathbb{R}^{d\times(d-r_{k,i})}$; $\bm{V}_{k}^{(i)}\in\mathcal{O}^{s_{k,i}}$ with $\bm{V}_{k,1}^{(i)}\in\mathbb{R}^{s_{k,i}\times r_{k,i}}$ and $\bm{V}_{k,2}^{(i)}\in\mathbb{R}^{s_{k,i}\times(s_{k,i}-r_{k,i})}$. This, together with $\bm{Q}_{i}\bm{Q}_{i}^{T}\bm{Z}_{k}^{(i)}=\bm{Z}_{k}^{(i)}$, implies for all $k\in[K]$ and $i\in[p]$,
$\bm{Q}_{i}\bm{Q}_{i}^{T}\bm{U}_{k,1}^{(i)}=\bm{U}_{k,1}^{(i)}.$ (49)
According to (5), (22), (44), and (45), we have for all $k\in[K]$ and $i\in[p]$,
$\alpha\bm{X}^{-1}\bm{Z}_{k}^{(i)}-\alpha\left(\bm{I}+\alpha_{k}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\right)^{-1}\bm{Z}_{k}^{(i)}=\lambda\bm{Z}_{k}^{(i)}.$ (50)
Substituting the block forms of $\bm{U}_{k}^{(i)}$ and $\bm{\Sigma}_{k}^{(i)}$ in (48) into the above equation and rearranging the terms, we obtain for all $k\in[K]$ and $i\in[p]$,
$\bm{X}^{-1}\bm{U}_{k,1}^{(i)}-\bm{U}_{k,1}^{(i)}\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{(i)^{2}}\right)^{-1}=\frac{\lambda}{\alpha}\bm{U}_{k,1}^{(i)}.$
Using $\bm{X}=\bm{I}+\alpha\bm{Z}\bm{Z}^{T}$ and rearranging the terms, we have for all $k\in[K]$ and $i\in[p]$,
$\bm{U}_{k,1}^{(i)}\left(\left(1-\frac{\lambda}{\alpha}\right)\bm{I}-\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{(i)^{2}}\right)^{-1}\right)=\alpha\bm{Z}\bm{Z}^{T}\bm{U}_{k,1}^{(i)}\left(\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{(i)^{2}}\right)^{-1}+\frac{\lambda}{\alpha}\bm{I}\right).$ (51)
Since $\bm{Z}\bm{Z}^{T}=\sum_{i=1}^{p}\lambda_{i}\bm{Q}_{i}\bm{Q}_{i}^{T}$, $\bm{Q}_{i}^{T}\bm{Q}_{j}=\bm{0}$ for all $i\neq j$, and (49), we have for all $k\in[K]$ and $i\in[p]$,
$\bm{U}_{k,1}^{(i)}\left(\left(1-\frac{\lambda}{\alpha}\right)\bm{I}-\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{(i)^{2}}\right)^{-1}\right)=\alpha\lambda_{i}\bm{U}_{k,1}^{(i)}\left(\left(\bm{I}+\alpha_{k}\tilde{\bm{\Sigma}}_{k}^{(i)^{2}}\right)^{-1}+\frac{\lambda}{\alpha}\bm{I}\right).$
Rearranging the terms in the above equation, we obtain for each $k\in[K]$ and $i\in[p]$,
$\tilde{\bm{\Sigma}}_{k}^{(i)}=\beta_{i}\bm{I},\ \text{where}\ \beta_{i}:=\frac{1}{\sqrt{\alpha K}}\sqrt{\frac{1+\alpha\lambda_{i}}{1-\lambda/\alpha-\lambda\lambda_{i}}-1}.$ (52)
Substituting this back to (51) yields for each $k\in[K]$ and $i\in[p]$,
$\lambda_{i}\bm{U}_{k,1}^{(i)}=\bm{Z}\bm{Z}^{T}\bm{U}_{k,1}^{(i)}.$
This, together with (48) and (52), yields $\lambda_{i}\bm{Z}_{k}^{(i)}=\bm{Z}\bm{Z}^{T}\bm{Z}_{k}^{(i)}$ for all $k\in[K]$ and $i\in[p]$. Using this and $\bm{Z}_{k}^{(i)^{T}}\bm{Z}_{k}^{(j)}=\bm{0}$ for all $i\neq j$, we have for all $i\in[p]$ and $k\in[K]$,
$\lambda_{i}\bm{Z}_{k}^{(i)}=\sum_{l=1}^{K}\bm{Z}_{l}^{(i)}\bm{Z}_{l}^{(i)^{T}}\bm{Z}_{k}^{(i)}.$
It follows from this and (46) that
$\lambda_{i}\bm{Z}^{(i)}=\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}\bm{Z}^{(i)}.$ (53)
Since there exists $k\neq l\in[K]$ such that $\bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0}$, we can assume without loss of generality that $\bm{Z}_{1}^{T}\bm{Z}_{2}\neq\bm{0}$. Then, there exist $i_{1}\in[m_{1}]$ and $i_{2}\in[m_{2}]$ such that $\bm{z}_{1,i_{1}}^{T}\bm{z}_{2,i_{2}}\neq 0$. This, together with $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$ for all $i\neq j$, implies that there exists $i^{*}\in[p]$ such that $\bm{z}_{1,i_{1}},\bm{z}_{2,i_{2}}$ are both columns of $\bm{Z}^{(i^{*})}$. Without loss of generality, suppose that $\bm{z}_{1,i_{1}}$ and $\bm{z}_{2,i_{2}}$ are the $u$-th and $v$-th columns of $\bm{Z}^{(i^{*})}$, respectively. Therefore, we have $\bm{z}_{u}^{(i^{*})^{T}}\bm{z}_{v}^{(i^{*})}\neq 0$. Using this, $\bm{z}_{u}^{(i^{*})^{T}}\bm{z}_{v}^{(i^{*})}\neq 0$, and (53), we have
$\lambda_{i^{*}}\bm{z}_{u}^{(i^{*})}=\bm{Z}^{(i^{*})}\bm{Z}^{(i^{*})^{T}}\bm{z}_{u}^{(i^{*})}.$ (54)
This is equivalent to
$\sum_{j\neq u}^{s_{i^{*}}}\bm{z}_{j}^{(i^{*})^{T}}\bm{z}_{u}^{(i^{*})}\bm{z}_{j}^{(i^{*})}+\left(\|\bm{z}_{u}^{(i^{*})}\|^{2}-\lambda\right)\bm{z}_{u}^{(i^{*})}=\bm{0}$ (55)
This, together with $\bm{z}_{u}^{(i^{*})^{T}}\bm{z}_{v}^{(i^{*})}\neq 0$, implies that the columns of $\bm{Z}^{(i^{*})}$ are linearly dependent. By letting $t_{i^{*}}=\mathrm{rank}(\bm{Z}^{(i^{*})})$, we have $t_{i^{*}}<s_{i^{*}}$ due to linear dependence of columns of $\bm{Z}^{(i^{*})}$. Then, let $\bm{Z}^{(i^{*})}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be an SVD of $\bm{Z}^{(i^{*})}$, where $\bm{U}\in\mathcal{O}^{d\times t_{i^{*}}}$, $\bm{\Sigma}\in\mathbb{R}^{t_{i^{*}}\times t_{i^{*}}}$, and $\bm{V}\in\mathcal{O}^{s_{i^{*}}\times t_{i^{*}}}$. Substituting this into (53) yields $\lambda_{i^{*}}\bm{\Sigma}=\bm{\Sigma}^{3}$, which implies $\bm{\Sigma}=\sqrt{\lambda_{i^{*}}}\bm{I}$ and
$\bm{Z}^{(i^{*})}=\sqrt{\lambda_{i^{*}}}\bm{U}\bm{V}^{T}.$ (56)
Construct an ascent direction.
For ease of exposition, we simply write $i^{*}$ as $i$ from now on. According to (46) and (54), we have
$\sum_{k=1}^{K}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\bm{z}_{u}^{(i)}=\lambda_{i}\bm{z}_{u}^{(i)}.$ (57)
Recall that $\bm{z}_{u}^{(i)}$ and $\bm{z}_{v}^{(i)}$ are a column of $\bm{Z}_{1}$ and $\bm{Z}_{2}$, respectively. Without loss of generality, suppose that $\bm{z}_{u}^{(i)}$ are the first column of $\bm{Z}_{1}^{(i)}$, i.e., $\bm{z}_{u}^{(i)}=\bm{Z}_{1}^{(i)}\bm{e}_{1}$. Then, let $\bm{c}=\left(\bm{c}_{1}\ \dots\ \bm{c}_{K}\right)\in\mathbb{R}^{s_{i}}$ with $\bm{c}_{1}=\bm{Z}_{1}^{(i)^{T}}\bm{z}_{u}^{(i)}-\lambda_{i}\bm{e}_{1}$ and $\bm{c}_{k}=\bm{Z}_{k}^{(i)^{T}}\bm{z}_{u}^{(i)}$ for all $k\neq 1$. This, together with $\bm{z}_{u}^{(i)^{T}}\bm{z}_{v}^{(i)}\neq 0$ and (57), implies $\bm{c}_{2}\neq\bm{0}$ and $\bm{Z}^{(i)}\bm{c}=\bm{0}$. Now, we set $\bm{q}_{k}:=\bm{V}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}\bm{c}_{k}$ for each $k\in[K]$ and $\bm{q}:=\left(\bm{q}_{1}\ \dots\ \bm{q}_{K}\right)$. According to $\bm{Z}_{k}^{(i)}=\beta_{i}\bm{U}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}$ by (48) and (52), we have for all $k\neq 1$,
$\bm{q}_{k}=\bm{V}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}\bm{Z}_{k}^{(i)^{T}}\bm{z}_{u}^{(i)}=\bm{Z}_{k}^{(i)^{T}}\bm{z}_{u}^{(i)}=\bm{c}_{k}.$
Moreover, using $\bm{Z}_{k}^{(i)}=\beta_{i}\bm{U}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}$ by (48) and (52), we have
$\bm{Z}^{(i)}\bm{q}=\sum_{k=1}^{K}\bm{Z}_{k}^{(i)}\bm{q}_{k}=\beta_{i}\sum_{k=1}^{K}\bm{U}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}\bm{c}_{k}=\sum_{k=1}^{K}\bm{Z}_{k}^{(i)}\bm{c}_{k}=\bm{Z}^{(i)}\bm{c}=\bm{0}$ (58)
and
$\|\bm{Z}_{k}^{(i)}\bm{q}_{k}\|=\beta_{i}\|\bm{V}_{k,1}^{(i)^{T}}\bm{c}_{k}\|=\beta_{i}\|\bm{q}_{k}\|.$ (59)
Let $\bm{u}=\bm{U}\bm{a}$, where $\bm{U}$ is given in (56) and $\bm{a}\in\mathbb{R}^{t_{i}}$ is chosen such that $\bm{a}\in\mathrm{span}(\bm{U}^{T}\bm{U}_{k,2}^{(i)})$ and $\|\bm{a}\|=1$. We construct $\bm{D}=\left[\bm{D}^{(1)}\ \dots\ \bm{D}^{(p)}\right]$ with $\bm{D}^{(i)}=\bm{u}\bm{q}^{T}$ and $\bm{D}^{(j)}=\bm{0}$ for all $j\neq i$.
Compute the bilinear form of Hessian.
According to the construction of $\bm{D}$ and (58), we compute $\bm{Z}\bm{D}^{T}=\bm{Z}^{(i)}\bm{D}^{(i)^{T}}=\bm{Z}^{(i)}\bm{q}\bm{u}^{T}=\bm{0}$. This, together with (23) and (47), yields
$\nabla^{2}R(\bm{Z})\left[\bm{D},\bm{D}\right]=\alpha\bm{a}^{T}\bm{U}^{T}\left(\bm{I}+\alpha\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}\right)^{-1}\bm{U}\bm{a}\|\bm{q}\|^{2}=\frac{\alpha}{\alpha\lambda_{i}+1}\|\bm{q}\|^{2},$
where the last equality is due to (56). With abuse of notation, let
$R_{c}\left(\bm{Z}_{k}^{(i)}\right)=\frac{m_{k}}{2m}\sum_{i=1}^{p}\log\det\left(\bm{I}_{n}+\alpha_{k}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\right).$
Since $\bm{Z}_{k}^{(i)}=\beta_{i}\bm{U}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}$ and $\bm{D}_{k}^{(i)}=\bm{u}\bm{q}_{k}^{T}$, we compute for each $k\in[K]$,
$\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$ $=\alpha\|\bm{q}_{k}\|^{2}\bm{u}^{T}\bm{X}_{k}^{(i)}\bm{u}-\alpha\alpha_{k}\left(\bm{u}^{T}\bm{X}_{k}^{(i)}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}-\alpha\alpha_{k}\left(\bm{u}^{T}\bm{X}_{k}^{(i)}\bm{u}\bm{q}_{k}^{T}\bm{Z}_{k}^{(i)^{T}}\bm{X}_{k}^{(i)}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)$
$=\alpha\|\bm{q}_{k}\|^{2}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}$
$\quad-\alpha\alpha_{k}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)\frac{\beta_{i}^{2}\|\bm{q}_{k}\|^{2}}{\alpha_{k}\beta_{i}^{2}+1},$
where $\bm{X}_{k}^{(i)}=\left(\bm{I}+\alpha_{k}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\right)^{-1}$, the second equality follows from
$\bm{X}_{k}^{(i)}=\left(\bm{I}+\alpha_{k}\beta_{i}^{2}\bm{U}_{k,1}^{(i)}\bm{U}_{k,1}^{(i)^{T}}\right)^{-1}=\bm{I}-{\alpha_{k}\beta_{i}^{2}}\bm{U}_{k,1}^{(i)}\bm{U}_{k,1}^{(i)^{T}}/({\alpha_{k}\beta_{i}^{2}+1})$
due to (26) in Lemma 2, $\bm{Z}_{k}^{(i)}=\beta_{i}\bm{U}_{k,1}^{(i)}\bm{V}_{k,1}^{(i)^{T}}$, and (59). Summing up the above equality for all $k\in[K]$ with $\alpha_{k}=K\alpha$ for all $k\in[K]$ yields
$\quad\sum_{k=1}^{K}\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$
$=\alpha\left(1-\frac{\alpha_{k}\beta_{i}^{2}}{\alpha_{k}\beta_{i}^{2}+1}\right)\|\bm{q}\|^{2}-\frac{\alpha\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
$\quad-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\frac{\alpha\alpha_{k}^{2}\beta_{i}^{2}}{\left(\alpha_{k}\beta_{i}^{2}+1\right)^{2}}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
$=\left(\frac{\alpha}{1+\alpha\lambda_{i}}-\lambda\right)\|\bm{q}\|^{2}-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right),$
where the second equality follows from the definition of $\beta_{i}$ in (52). Finally, we compute
$\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]$ $=\nabla^{2}R(\bm{Z})[\bm{D},\bm{D}]-\sum_{j=1}^{K}\nabla^{2}R_{c}(\bm{z}_{j}^{(i)})[\bm{d}_{j}^{(i)},\bm{d}_{j}^{(i)}]-\lambda\|\bm{q}\|^{2}$
$=\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)>0,$
where the inequality is due to $\|\bm{q}_{2}\|=\|\bm{c}_{2}\|\neq 0$ and
$\|\bm{Z}_{2}^{(i)^{T}}\bm{u}\|^{2}=\beta_{2}\|\bm{U}_{2,1}^{(i)^{T}}\bm{U}\bm{a}\|\neq 0$
due to $\bm{a}\in\mathrm{span}(\bm{U}^{T}\bm{U}_{k,2}^{(i)})$. ∎
Given a matrix $\bm{Z}\in\mathbb{R}^{d\times m}$, let $\bm{Z}\bm{Z}^{T}=\bm{Q}\bm{\Lambda}\bm{Q}^{T}$ be an eigenvalue decomposition of $\bm{Z}\bm{Z}^{T}\in\mathbb{S}_{+}^{d}$, where $\bm{Q}\in\mathcal{O}^{d\times r}$ and $\bm{\Lambda}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with diagonal entries being positive eigenvalues of $\bm{Z}\bm{Z}^{T}$. Suppose that $\bm{Z}\bm{Z}^{T}$ has $p$ distinct positive eigenvalues, where $1\leq p\leq r$. Let $\lambda_{1}>\dots>\lambda_{p}>0$ be its distinct eigenvalue values with the corresponding multiplicities being $h_{1},\dots,h_{p}\in\mathbb{N}_{+}$, respectively. Obviously, we have $\sum_{i=1}^{p}h_{i}=r$. Therefore, we write
$\bm{\Lambda}=\mathrm{BlkDiag}\left(\lambda_{1}\bm{I}_{h_{1}},\dots,\lambda_{p}\bm{I}_{h_{p}}\right),\ \bm{Q}=\begin{bmatrix}\bm{Q}_{1},\dots,\bm{Q}_{p}\end{bmatrix},$
where $\bm{Q}_{i}\in\mathcal{O}^{d\times h_{i}}$ for all $i\in[p]$.
Appendix D Proofs in Section 3
D.1 Proof of Theorem 1
Proof of Theorem 1.
(i) Suppose that each block of $\bm{Z}$ satisfying (7), (a), (b), and (c). It directly follows from (i) in Proposition 3 that $\bm{Z}$ is a local maximizer. Conversely, suppose that $\bm{Z}$ is a local maximizer. According to (17), (18), (19), (ii) in Proposition 3 and Proposition 4, if $\bm{Z}\in\mathcal{Z}^{c}\cup\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point. This, together with $\mathcal{X}=\mathcal{Z}^{c}\cup\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$ and the fact that $\bm{Z}$ is a local maximizer, implies that $\bm{Z}\in\mathcal{Z}_{1}$. Using this, (19), and Proposition 2 yields that $\bm{Z}$ satisfying (7), (a), (b), and (c).
(ii) According to (i) in Theorem 1, suppose that the $k$-th block of a local maximizer $\bm{Z}$ admits the decomposition in (7) satisfying (a), (b), and (c) for all $k\in[K]$. This, together with (C.1) in the proof of Proposition 3, yields that
$F(\bm{Z})$ $=\frac{1}{2}\sum_{k=1}^{K}\sum_{i=1}^{r_{k}}\left(\log\left(1+\alpha\overline{\sigma}_{k}^{2}\right)-\frac{m_{k}}{m}\log\left(1+\alpha_{k}\overline{\sigma}_{k}^{2}\right)-\lambda\overline{\sigma}_{k}^{2}\right),$ (60)
where $\overline{\sigma}_{k}$ is defined in (16) for each $k\in[K]$. Then, we define a function $g:\mathbb{N}_{+}\times\mathbb{R}\rightarrow\mathbb{R}$ as
$g(n,x):=\log(1+\alpha x)-\frac{n}{m}\log\left(1+\frac{dx}{n\varepsilon^{2}}\right)-\lambda x.$
One can verify that for all $n_{1}\geq n_{2}$, we have $g(n_{1},x)\leq g(n_{2},x)$ for each $x$. Therefore, we have for all $m_{k}\leq m_{l}$,
$g(m_{l},\overline{\sigma}_{l}^{2})\leq g(m_{k},\overline{\sigma}_{l}^{2})\leq g(m_{k},\overline{\sigma}_{k}^{2}),$
where the second inequality follows from $\overline{\sigma}_{k}^{2}$ is the maximizer of the function $g(m_{k},x)=h_{k}(x)$ according to (42). This, together with (60), yields that $\bm{Z}$ is a global maximizer if and only if $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$ and for all $k\neq l$ satisfying $m_{k}<m_{l}$ and $r_{l}>0$, we have $r_{k}=\min\{m_{k},d\}$. ∎
D.2 Proof of Proposition 1
To prove Proposition 1, we first need to characterize the global optimal solution set of Problem (4).
Proposition 5.
Suppose that $m_{1}=\dots=m_{K}$ and (8) holds. It holds that $\bm{Z}=[\bm{Z}_{1},\dots,\bm{Z}_{K}]\in\mathbb{R}^{d\times m}$ with $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$ for each $k\in[K]$ is a global solution of Problem (4) if and only if for each $k\in[K]$,
$\bm{Z}_{k}=\frac{m}{\min\{m,d\}}\bm{U}_{k}\bm{V}_{k}^{T},$ (61)
where $r_{k}=\min\{m,d\}/K$ for all $k\in[K]$, $\bm{U}_{k}\in\mathcal{O}^{d\times r_{k}}$ with $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $l\neq k$, and $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$ for all $k\in[K]$.
Proof.
According to Lemma 4, we have
$R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)$ $\leq\frac{1}{2}\sum_{k=1}^{K}\log\det\left(\bm{I}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\sum_{k=1}^{K}\frac{m_{k}}{2m}\log\det\left(\bm{I}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)$
$=:H(\bm{Z}),$ (62)
where the inequality becomes equality if and only if $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$. To simplify our development, let $r_{k}:=\mathrm{rank}(\bm{Z}_{k})$ denote the rank of $\bm{Z}_{k}\in\mathbb{R}^{d\times m_{k}}$, where $r_{k}\leq\min\{d,m_{k}\}$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq\min\{d,m\}$, and
$h_{k}(\bm{Z}_{k}):=\frac{1}{2}\log\det\left(\bm{I}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right),\ \forall k\in[K].$ (63)
Moreover, let
$\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}\bm{Q}_{k}^{T}=\begin{bmatrix}\bm{P}_{k,1}&\bm{P}_{k,2}\end{bmatrix}\begin{bmatrix}\tilde{\bm{\Sigma}}_{k}&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}\begin{bmatrix}\bm{Q}_{k,1}^{T}\\ \bm{Q}_{k,2}^{T}\end{bmatrix}$
be a singular value decomposition (SVD) of $\bm{Z}_{k}$, where $\tilde{\bm{\Sigma}}_{k}=\mathrm{diag}(\sigma_{k,1},\dots,\sigma_{k,r_{k}})$ with $\sigma_{k,1}\geq\dots\geq\sigma_{k,r_{k}}>0$ being positive singular values of $\bm{Z}_{k}$, $\bm{P}_{k}\in\mathcal{O}^{d}$ with $\bm{P}_{k,1}\in\mathbb{R}^{d\times r_{k}}$ and $\bm{P}_{k,2}\in\mathbb{R}^{d\times(d-r_{k})}$, and $\bm{Q}_{k}\in\mathcal{O}^{m_{k}}$ with $\bm{Q}_{k,1}\in\mathbb{R}^{m_{k}\times r_{k}}$ and $\bm{Q}_{k,2}\in\mathbb{R}^{m_{k}\times(m_{k}-r_{k})}$. Substituting this SVD into (63), together with $\|\bm{Z}_{k}\|_{F}^{2}=m_{k}$, yields that to maximize $H(\bm{Z})$, it suffices to study for each $k\in[K]$,
$\max_{\sigma_{k,1},\dots,\sigma_{k,r_{k}}}\sum_{i=1}^{r_{k}}\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\sum_{i=1}^{r_{k}}\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\qquad\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}\sigma_{k,i}^{2}=m_{k}.$
To simplify our development, let $x_{i}:=\sigma_{k,i}^{2}\geq 0$ for each $i\in[r_{k}]$. This, together with $m_{k}=m/K$, implies that it suffices to study
$\max_{x_{1},\dots,x_{r_{k}}}g(\bm{x})$ $:=\sum_{i=1}^{r_{k}}\log\left(1+\alpha x_{i}\right)-\sum_{i=1}^{r_{k}}\frac{1}{K}\log\left(1+\alpha_{k}x_{i}\right)\qquad$
$\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}x_{i}=\frac{m}{K},\ \ x_{i}\geq 0,\ \forall i\in[r_{k}].$ (64)
This, together with Lemma 6 and (11), yields that the optimal solution for each $k\in[K]$ is
$x_{i}^{*}=\frac{m}{r_{k}K},\ \forall i\in[r_{k}].$ (65)
(i) Suppose that $m\leq d$. Then, we have $r_{k}\leq m/K$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq m$. This, together with (65) and Lemma 7, implies that $r_{k}=m/K$, and thus $x_{i}^{*}=1$ for all $i\in[r_{k}]$ and $k\in[k]$.
(ii) Suppose that $m>d$. Then, we have $r_{k}\leq\min\{d,m/K\}$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq d$. To compute the optimal function value, we consider the following problem:
$\max_{r_{1},\dots,r_{K}\in\mathbb{Z}}$ $\sum_{k=1}^{K}r_{k}\left(\log\left(1+\frac{m\alpha}{r_{k}K}\right)-\frac{1}{K}\log\left(1+\frac{m\alpha}{r_{k}}\right)\right)$
$\mathrm{s.t.}\ \sum_{k=1}^{K}r_{k}=d,\ r_{k}\leq\min\left\{d,\frac{m}{K}\right\},\ \forall k\in[K].$
Now, we study the following function:
$\phi(x):=x\left(\log\left(1+\frac{m\alpha}{Kx}\right)-\frac{1}{K}\log\left(1+\frac{m\alpha}{x}\right)\right),\ \text{where}\ x\in\left[1,\frac{m}{K}\right]$
We compute
$\phi^{\prime}(x)$ $=\log\left(1+\frac{m\alpha}{Kx}\right)-\frac{1}{K}\log\left(1+\frac{m\alpha}{x}\right)-\frac{m\alpha}{Kx+m\alpha}+\frac{m\alpha}{K(x+m\alpha)},$
$\phi^{\prime\prime}(x)$ $=-\frac{m\alpha/x}{Kx+m\alpha}+\frac{m\alpha/x}{K(x+m\alpha)}+\frac{Km\alpha}{(Kx+m\alpha)^{2}}-\frac{m\alpha}{K(x+m\alpha)^{2}}$
$=-\frac{m^{2}\alpha^{2}/x}{(Kx+m\alpha)^{2}}+\frac{m^{2}\alpha^{2}/x}{K(x+m\alpha)^{2}}.$
Since $x\in[1,m/K]$, we have $Kx^{2}\leq m^{2}/K\leq m^{2}\alpha^{2}$ when $\alpha\geq 1/\sqrt{K}$, and thus $\phi^{\prime\prime}(x)\leq 0$. Therefore, $\phi(x)$ is a concave function for all $x\in[1,m/K]$. Then, applying the Jensen inequality yields
$\sum_{k=1}^{K}\frac{1}{K}f(r_{k})\leq f\left(\sum_{k=1}^{K}r_{k}\right),$
where the inequality becomes equality if and only if $r_{1}=\dots=r_{k}=d/K$. This, together with (65), yields $x_{i}^{*}=m/d$.
According to (i) and (ii), we have $x_{i}^{*}=m/\min\{m,d\}$ and $r_{k}=\min\{m,d\}/K$. Therefore, we have $\bm{Z}_{k}=m/\min\{m,d\}\bm{P}_{k,1}\bm{Q}_{k,1}^{T}$, where $\bm{P}_{k,1}\in\mathcal{O}^{d\times r_{k}}$ and $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$ for each $k\in[K]$. Then, we complete the proof.
Based on the above proposition, we are ready to prove Proposition 1.
Proof of Proposition 1.
Let $\hat{\bm{Z}}=[\bm{Z}_{1},\dots,\bm{Z}_{k}]$ denote the optimal solution of Problem (4). According to Proposition 5, it suffices to study the following two cases.
(i) Suppose that $m<d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=\bm{U}_{k}\bm{V}_{k}^{T}$ and $\hat{r}_{k}=m/K$ for each $k\in[K]$. Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (10), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=1$ and $\sum_{k=1}^{K}r_{k}=m$. Since $r_{k}\leq m_{k}$ for each $k\in[K]$, we have $r_{k}=m/K$ for each $k\in[K]$. Therefore, Problem (4) and Problem (5) have the same global solution set.
(ii) Suppose that $m\geq d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=m\bm{U}_{k}\bm{V}_{k}^{T}/d$ and $\hat{r}_{k}=d/K$ for each $k\in[K]$. Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (9), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=m/d$ and $\sum_{k=1}^{K}r_{k}=d$. Therefore, the global solution set of Problem (4) is a subset of that of Problem (5).
Lemma 6.
Suppose that $m,K$ are integers such that $m/K$ is a positive integer, $r\leq m/K$ is an integer, and $\alpha>0$ is a constant. Consider the following optimization problem
$\min_{\bm{x}\in\mathbb{R}^{r}}\sum_{i=1}^{r}-\log\left(1+\alpha x_{i}\right)+\sum_{i=1}^{r}\frac{1}{K}\log(1+K\alpha x_{i})\quad\mathrm{s.t.}\ \sum_{i=1}^{r}x_{i}=\frac{m}{K},\ x_{i}\geq 0,\ \forall i\in[r].$ (66)
If
$\alpha\geq 6K^{\frac{1}{K-1}}\exp(1)\left(1+\frac{1}{\sqrt{K}}\right)^{\frac{m}{K-1}},$ (67)
the optimal solution is
$x_{i}^{*}=\frac{m}{rK},\ \forall i\in[r].$ (68)
Proof.
If $r=1$, it is trivial to see that (68) is the optimal solution. Therefore, it suffices to study $r\geq 2$. To simplify our development, let $f(x):=-\log(1+\alpha x)+\log(1+K\alpha x)/K$ and $F(\bm{x}):=\sum_{i=1}^{r}f(x_{i})$. Then, one can verify that for all $x\geq 0$,
$f^{\prime}(x)=-\frac{\alpha}{1+\alpha x}+\frac{\alpha}{1+K\alpha x}<0,\quad f^{\prime\prime}(x)=\frac{\alpha^{2}}{(1+\alpha x)^{2}}-\frac{K\alpha^{2}}{(1+K\alpha x)^{2}}.$ (69)
Introducing dual variables $\lambda$ associated with the constraint $\sum_{i=1}^{r}x_{i}=m/K$ and $\mu_{i}$ associated with the constraint $x_{i}\geq 0$ for each $i\in[r]$, we write the Lagrangian as follows
$\mathcal{L}(\bm{x};\lambda,\bm{\mu})=\sum_{i=1}^{r}f(x_{i})+\lambda\left(\sum_{i=1}^{r}x_{i}-\frac{m}{K}\right)-\sum_{i=1}^{r}\mu_{i}x_{i}.$ (70)
Then, we write the KKT system as follows:
$-\frac{\alpha}{1+\alpha x_{i}}+\frac{\alpha}{1+K\alpha x_{i}}+\lambda-\mu_{i}=0,\ x_{i}\mu_{i}=0,\ x_{i}\geq 0,\ \mu_{i}\geq 0,\ \forall i\in[r],\ \sum_{i=1}^{r}x_{i}=\frac{m}{K}.$ (71)
Now, let $\mathcal{S}:=\{i\in[r]:x_{i}>0\}$ denote the support set of a KKT point $\bm{x}\in\mathbb{R}^{r}$ and $s:=|\mathcal{S}|$ denote the cardinality of the support set, where $1\leq s\leq r$. This, together with (71), implies that for each $i\in\mathcal{S}$,
$-\frac{\alpha}{1+\alpha x_{i}}+\frac{\alpha}{1+K\alpha x_{i}}+\lambda=0,\ \sum_{i\in\mathcal{S}}x_{i}=\frac{m}{K}.$ (72)
This is equivalent to the following quadratic equation:
$K\lambda\alpha x_{i}^{2}-\left((K-1)\alpha-(K+1)\lambda\right)x_{i}+\frac{\lambda}{\alpha}=0.$ (73)
We compute
$\Delta=\eta^{2}-4K\lambda^{2},\text{where}\ \eta:=(K-1)\alpha-(K+1)\lambda.$ (74)
Note that for all $i\in\mathcal{S}$, we have $x_{i}>0$, and thus $\mu_{i}=0$. This, together with $K\geq 2$ and the first equation in (72), implies $\lambda>0$. Consequently, the quadratic equation (73) has a positive root if and only if $\eta\geq 0$ and $\Delta\geq 0$. This implies
$0<\lambda\leq\frac{\sqrt{K}-1}{\sqrt{K}+1}\alpha.$ (75)
Then, the solution of Problem (73) is $x_{i}\in\{\overline{x},\underline{x}\}$ for each $i\in\mathcal{S}$, where
$\overline{x}=\frac{\eta+\sqrt{\Delta}}{2K\lambda\alpha},\quad\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}.$ (76)
Now, we discuss the KKT points that could potentially be optimal solutions. Let $\bm{x}\in\mathbb{R}^{r}$ be a KKT point satisfying $x_{i}\in\{\overline{x},\underline{x}\}$ for each $i\in\mathcal{S}$, where $s\in\{1,2,\dots,r\}$. In particular, when $s=1$, we have $x_{i}=m/K$ for all $i\in\mathcal{S}$. In the following, we consider $s\in\{2,\dots,r\}$.
Case 1. Suppose that $x_{i}=x_{j}$ for all $i,j\in\mathcal{S}$. This, together with $\sum_{i\in\mathcal{S}}x_{i}=m/K$ and $x_{i}\in\{\overline{x},\underline{x}\}$ for each $i\in\mathcal{S}$, yields
$x_{i}=\frac{m}{sK},\ \forall i\in\mathcal{S}.$ (77)
Case 2. Suppose that there exists $i\neq j\in\mathcal{S}$ such that $x_{i}\neq x_{j}$. This, together with $x_{i}\in\{\overline{x},\underline{x}\}$, implies $\overline{x}>\underline{x}$. According to (69), we have $f^{\prime\prime}(x)=0$ at $\hat{x}=1/(\alpha\sqrt{K})$. Then, we obtain that $f^{\prime}(x)$ is strictly decreasing in $[0,\hat{x}]$ and strictly increasing in $[\hat{x},\infty]$. Then, one can further verify that $\underline{x}<\hat{x}<\overline{x}$. This, together with (69), implies
$f^{\prime\prime}(\underline{x})<0,\ f^{\prime\prime}(\overline{x})>0.$ (78)
For ease of exposition, let $l(\bm{x})=|\{i\in\mathcal{S}:x_{i}=\underline{x}\}|$ be the number of entries of $\bm{x}$ that equal to $\underline{x}$. Then, we claim that any optimal solution $\bm{x}^{\star}$ satisfies $l(\bm{x}^{\star})\leq 1$. Now, we prove this claim by contradiction. Without loss of generality, we assume that $x_{i}^{\star}=\overline{x}$ for all $i=1,\dots,r-l$ and $x_{i}^{\star}=\underline{x}$ for all $i=r-l+1,\dots,r$ with $l\geq 2$. This, together with (78) and $l\geq 2$, yields
$f^{\prime\prime}(x_{r-1}^{\star})<0,\quad f^{\prime\prime}(x_{r}^{\star})<0.$ (79)
Using the second-order necessary condition for constraint optimization problems (see, e.g., [35, Theorem 12.5]) and $x_{i}^{\star}\geq 0$ for all $i\in[r]$, we obtain
$\sum_{i=1}^{r}f^{\prime\prime}(x_{i}^{\star})v_{i}^{2}\geq 0,\ \forall\bm{v}\in\mathbb{R}^{r}\ \text{s.t.}\ \sum_{i=1}^{r}v_{i}=0.$ (80)
Then, we take $\bm{v}\in\mathbb{R}^{r}$ such that $v_{1}=\dots=v_{r-2}=0$ and $v_{r-1}=-v_{r}\neq 0$. Substituting this into (80) yields
$f^{\prime\prime}(x_{r-1}^{\star})+f^{\prime\prime}(x_{r}^{\star})\geq 0,$
which contradicts (79). Therefore, $\bm{x}^{\star}$ cannot be an optimal solution. Then, we prove the claim. In this case, we can write the KKT point that could be an optimal solution as follows: There exists $i\in\mathcal{S}$ such that
$x_{i}=\underline{x},\ x_{j}=\overline{x},\ \forall j\neq i.$ (81)
Now, we compute the function values for the above two cases, i.e., (77) and (81), and compare them to determine which one is the optimal solution. To simplify our analysis, let $h(x):=f(x)/x$. For (77) in Case 1, we compute
$F_{1}=sf\left(\frac{m}{sK}\right)=\frac{m}{K}h\left(\frac{m}{sK}\right).$ (82)
For (81) in Case 2, we compute
$F_{2}=(s-1)f\left(\overline{x}\right)+f\left(\underline{x}\right)\geq(s-1)f\left(\frac{m}{(s-1)K}\right)+f(\underline{x})=\frac{m}{K}h\left(\frac{m}{(s-1)K}\right)+f(\underline{x}),$ (83)
where the inequality is because $\overline{x}\leq{m}/\left({(s-1)K}\right)$ and $f(x)$ is strictly dereasing in $[0,\infty)$. For all $x\in\left[{m/(sK)},{m}/\left({(s-1)K}\right)\right]$, we compute
$h^{\prime}(x)$ $=\frac{f^{\prime}(x)x-f(x)}{x^{2}}=-\frac{1}{x}\left(\frac{\alpha}{1+\alpha x}-\frac{\alpha}{1+K\alpha x}\right)+\frac{1}{x^{2}}\left(\log(1+\alpha x)-\frac{1}{K}\log(1+K\alpha x)\right)$
$\geq-\frac{(K-1)\alpha^{2}}{(1+\alpha x)(1+K\alpha x)}+\frac{1}{x^{2}}\left(\log(1+\alpha)-\frac{1}{K}\log(1+\alpha K)\right)$
$\geq-\frac{K-1}{Kx^{2}}\left(1-\log(1+\alpha)+\frac{\log K}{K-1}\right)\geq-\frac{(K-1)Ks^{2}}{m^{2}}\left(1+\frac{\log K}{K-1}-\log(1+\alpha)\right),$ (84)
where the first inequality follows from $\log(1+\alpha x)-\log(1+K\alpha x)/K\geq\log(1+\alpha)-\log(1+\alpha K)/K$ due to $x\geq m/(sK)\geq 1$, the second inequality uses $\log(1+\alpha)-\log(1+\alpha K)/K=(K-1)\log(1+\alpha)/K+\log\left((1+\alpha)/(1+\alpha K)\right)/K\geq\left((K-1)\log(1+\alpha)-\log K\right)/K$, and the last inequality is because of $x\geq{m}/(sK)$. According to (67), we have
$0\displaystyle 1+\frac{\log K}{K-1}-\log(1+\alpha)=\log\left(\frac{{K}^{\frac{1}{K-1}}\exp(1)}{1+\alpha}\right)<$ 0 . (85)
Using the mean-value theorem, there exists $x\in\left({m/(sK)},{m}/\left({(s-1)K}\right)\right)$ such that
$h\left(\frac{m}{(s-1)K}\right)-h\left(\frac{m}{sK}\right)=h^{\prime}(x)\left(\frac{m}{(s-1)K}-\frac{m}{sK}\right)\geq\frac{(K-1)s}{m(s-1)}\left(\log(1+\alpha)-1-\frac{\log K}{K-1}\right),$ (86)
where the inequality follows from (D.2). Now, we are devoted to bounding $f(\underline{x})$. According to (74) and (75), we have
$\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}=\frac{4K\lambda^{2}}{2K\lambda\alpha(\eta+\sqrt{\Delta})}\leq\frac{2\lambda}{\alpha\eta}.$ (87)
This, together with the fact that $f(x)$ is decreasing in $(0,\infty)$, yields
$f(\underline{x})$ $\geq f\left(\frac{2\lambda}{\alpha\eta}\right)=-\log\left(1+\frac{2\lambda}{\eta}\right)+\frac{1}{K}\log\left(1+\frac{2K\lambda}{\eta}\right)\geq-\log\left(1+\frac{2\lambda}{\eta}\right)\geq-\log\left(1+\frac{1}{\sqrt{K}}\right),$
where the last inequality uses $\eta=(K-1)\alpha-(K+1)\lambda\geq 2\sqrt{K}\lambda$ due to $(K-1)\alpha\geq(\sqrt{K}+1)^{2}\lambda$ by (75). This, together with (82), (83), and (86), yields
$F_{2}-F_{1}$ $=\frac{m}{K}\left(h\left(\frac{m}{(s-1)K}\right)-h\left(\frac{m}{sK}\right)\right)+f(\underline{x})$
$\geq\frac{(K-1)s}{m(s-1)}\left(\log(1+\alpha)-1-\frac{\log K}{K-1}\right)-\log\left(1+\frac{1}{\sqrt{K}}\right)$
$\geq\frac{K-1}{m}\log\left(\frac{1+\alpha}{{K}^{\frac{1}{K-1}}\exp(1)}\right)-\log\left(1+\frac{1}{\sqrt{K}}\right)>0,$
where the last inequality follows from (67). This implies that the optimal solution takes the form of (77) for some $s\in[r]$. Consequently, the function value of (77) for each $s\in[r]$ is
$s\left(-\log\left(1+\frac{\alpha m}{sK}\right)+\frac{1}{K}\log\left(1+\frac{\alpha m}{s}\right)\right).$
This, together with Lemma 7, implies that when the optimal solution takes the form of (77) with $s=r$, Problem (66) achieves its global minimum. Then, we complete the proof. ∎
Lemma 7.
Consider the setting in Lemma 6 and the following function
$h(s):=s\left(\frac{1}{K}\log\left(1+\frac{m\alpha}{s}\right)-\log\left(1+\frac{m\alpha}{sK}\right)\right),$ (88)
where $s\in[1,r]$ and $\alpha$ satisfies (67). Then, $h(s)$ is decreasing in $s\in[1,r]$.
Proof.
For ease of exposition, let $\beta:=m\alpha$ and $x:=1/s\in[1/r,1]$. According to (67), we have
$\alpha\geq 6K^{\frac{1}{K-1}}\exp(1)\left(1+\frac{1}{\sqrt{K}}\right)^{\frac{2m}{K-1}}>1\geq\frac{r\sqrt{K}}{m}.$
This implies $\beta\geq r\sqrt{K}$. Then, we study
$h(s)=g(x)=\frac{1}{x}\left(\frac{1}{K}\log\left(1+\beta x\right)-\log\left(1+\frac{\beta x}{K}\right)\right).$ (89)
Note that showing $h(s)$ is decreasing in $s\in[1,r]$ is equivalent to proving $g(x)$ is increasing in $x\in[1/r,1]$. Now, we compute
$g^{\prime}(x)$ $=\frac{1}{x}\left(\frac{\beta}{K(1+\beta x)}-\frac{\beta}{K+\beta x}\right)-\frac{1}{x^{2}}\left(\frac{1}{K}\log\left(1+\beta x\right)-\log\left(1+\frac{\beta x}{K}\right)\right)=-\frac{1}{x^{2}}\phi(x),$ (90)
where
$\phi(x):=\frac{1}{K}\log\left(1+\beta x\right)-\log\left(1+\frac{\beta x}{K}\right)+\beta x\left(\frac{1}{K+\beta x}-\frac{1}{K(1+\beta x)}\right).$
Then, it suffices to show $\phi(x)\leq 0$ for all $x\in[K/m,1]$ due to $1/r\geq K/m$. Towards this goal, we compute
$\phi^{\prime}(x)$ $=\frac{\beta}{K(1+\beta x)}-\frac{\beta}{K+\beta x}+\beta\left(\frac{1}{K+\beta x}-\frac{1}{K(1+\beta x)}\right)+\beta x\left(\frac{-\beta}{(K+\beta x)^{2}}+\frac{\beta}{K(1+\beta x)^{2}}\right)$
$=-x\beta^{2}\frac{(K-1)(\beta^{2}x^{2}-K)}{(K+\beta x)^{2}K^{2}(1+\beta x)^{2}}\leq 0,$
where the inequality follows from $\beta^{2}x^{2}\geq\beta^{2}K^{2}/m^{2}\geq K$ due to $x\in[K/m,1]$ and $\beta=\alpha m>2m>m/\sqrt{K}$. Therefore, $\phi(x)$ is decreasing in $[K/m,1]$. Next, we compute
$\phi\left(\frac{K}{m}\right)$ $=\frac{1}{K}\log\left(1+\frac{\beta K}{m}\right)-\log\left(1+\frac{\beta}{m}\right)+\frac{\beta}{m}\left(\frac{1}{1+{\beta}/{m}}-\frac{1}{1+{\beta K}/{m}}\right)$
$=\frac{1}{K}\log\left(1+\alpha K\right)-\log\left(1+\alpha\right)+\alpha\left(\frac{1}{1+\alpha}-\frac{1}{1+\alpha K}\right)$
$\leq\frac{1}{K}\log\left(1+\alpha K\right)-\log\left(1+\alpha\right)+1\leq\frac{1}{2}\log\left(1+2\alpha\right)-\log\left(1+\alpha\right)+1\leq 0,$
where the second equality is due to $\beta=\alpha m$, the second inequality holds because $\log\left(1+\alpha K\right)/K$ is decreasing as $K\geq 2$ increases when $\alpha\geq 2$, and the last inequality follows from $\alpha\geq 15$ by (67). This, together with the fact that $\phi$ is decreasing in $[K/m,1]$, yields $\phi(x)\leq\phi\left({K}/{m}\right)\leq 0$. Using this and (90), we obtain $g^{\prime}(x)\geq 0$ in $[K/m,1]$. Therefore, $g(x)$ is increasing in $[K/m,1]$. Then, we complete the proof. ∎
D.3 Proof of and Theorem 2
Proof of Theorem 2.
According to (i) of Proposition 3, if $\bm{Z}$ is a critical point but not a local maximizer, we have $\bm{Z}\in\mathcal{Z}_{2}\cup\mathcal{Z}^{c}$. This, together with (ii) of Proposition 3 and Proposition 4, yields that $\bm{Z}$ is a strict saddle point. ∎
Appendix E Additional Experimental Setups and Results
In this section, we provide additional implementation details and experimental results under different parameter settings for Sections 5.1 and 5.2 in Sections E.1 and E.2, respectively.
E.1 Implementation Details and Additional Results in Section 5.1
Training setups.
In this subsection, we employ full-batch gradient descent (GD) for solving Problem (5). Here, we use the Gaussian distribution to randomly initialize GD. More precisely, we randomly generate an initial point $\bm{Z}^{(0)}$ whose entries are i.i.d. sampled from the standard normal distribution, i.e. $z_{ij}^{(0)}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. We fix the learning rate of GD as $10^{-1}$ in the training. We terminate the algorithm when the gradient norm at some iterate is less $10^{-5}$.
(a) Experiment 1
(b) Experiment 2
(c) Experiment 3
Figure 5: Heatmap of cosine similarity between pairwise features under different settings.
In addition to the results presented in Section 5.1, we perform additional experiments under different settings as follows. To support our theorems, we visualize the heatmap of learned features and plot the number and magnitude of singular values in each class. Unless specified otherwise, we use the training setups introduced above in the following experiments.
  • Experiment 1 on balanced data. In this experiment, we consider that the number of samples in each class is same and set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=m_{2}=m_{3}=m_{4}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(a) and the number and magnitude of singular values in each class in Figure 6.
  • Experiment 2 on data with more classes. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=8$, the number of samples in each class is $m_{1}=50,m_{2}=40,m_{3}=30,m_{4}=20,m_{5}=30,m_{6}=40,m_{7}=40,m_{8}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(b) and the number and magnitude of singular values in each class in Figure 7.
  • Experiment 3 on data where the dimension $d$ is larger than the number of samples $m$. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=300$, the number of classes $K=4$, the number of samples in each class is $m_{1}=50,m_{2}=50,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.01$, and the quantization error $\epsilon=5$. Note that in this experiment, we set the learning rate as $1$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(c) and the number and magnitude of singular values in each class in Figure 8.
According to the results in Figures 5(a), 5(b), 6, and 7 of Experiments 1 and 2, we observe that when the number of samples is larger than its dimension, i.e., $m\geq d$, the learned features via the MCR2 principle are within-class compressible and between-class discriminative in both balanced and unbalanced data sets. Moreover, the dimension of the space spanned by these features is maximized such that $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$. This directly supports Theorem 1. By comparing the function value returned by GD and that computed by the closed-form in Theorem 1, we found that GD with random initialization will always converge to a global maximizer of Problem (5) when the data is balanced, while it will always converge to a local maximizer of Problem (5) when data is unbalanced. This directly supports Theorem 2.
According to the results in Figures 5(c) and 8 of Experiment 3, we observe that when the number of samples is smaller than its dimension, i.e., $m\leq d$, the learned features via the MCR2 principle are orthogonal to each other and the dimension of each subspace is equal to the number of samples, i.e., $r_{k}=m_{k}$ for each $k\in[K]$. This exactly supports Theorem 1. Indeed, when $d\geq m$ and $r_{k}=m_{k}$ for each $k\in[K]$, it follows from Theorem 1 that $\bm{V}_{k}=\bm{I}$ and thus $\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}$ for each $k\in[K]$ for each local maximizer. Therefore, this also supports Theorem 2 as GD with random initialization converges to a local maximizer.
Figure 6: The number and magnitude of singular values in each subspace in Experiment 1. The blue dots are plotted based on the singular values by applying SVD to the solution returned by GD, and the red line is plotted according to the closed-form solution in (7). The number of singular values in each subspace is $25,24,24,27$, respectively.
Figure 7: The number and magnitude of singular values in each subspace in Experiment 2. The blue dots are plotted based on the singular values by applying SVD to the solution returned by GD, and the red line is plotted according to the closed-form solution in (7). The number of singular values in each subspace is $8,13,13,17,15,13,12,9$, respectively.
Figure 8: The number and magnitude of singular values in each subspace in Experiment 3. The blue dots are plotted based on the singular values by applying SVD to the solution returned by GD, and the red line is plotted according to the closed-form solution in (7). The number of singular values in each subspace is $50,50,40,60$, respectively.
MNIST: $m=1000,K=4$
MNIST: $m=2000,K=8$
CIFAR: $m=1000,K=4$
CIFAR: $m=2000,K=8$
Figure 9: Heatmap of cosine similarity among learned features by training deep networks on MNIST and CIFAR-10. We train network parameters by optimizing the regularized MCR2 objective (5) on $m$ samples split equally among $K$ classes of MNIST and CIFAR-10. In the figure, the darker pixels represent higher cosine similarity between features. In particular, when the $(i,j)$-th pixel is close to 00 (very light blue), the features $i$ and $j$ are approximately orthogonal.
E.2 Implementation Details and Additional Results in Section 5.2
Network architecture and training setups for MNIST.
In the experiments on MNIST, we employ a 4-layer multilayer perception (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 2048 and output dimension 32. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28]. We terminate the algorithm when it reaches 3000 epochs.
More experimental results on MNIST.
Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem 1.
Network architecture and training setups for CIFAR-10.
In the experiments on CIFAR10, we employ a 2-layer multilayer perceptron (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 4096 and output dimension 128. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28]. We terminate the algorithm when it reaches 4000 epochs.
More experimental results on CIFAR-10.
Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem 1.

### v2

@@
 1.1 Background and Motivation
-In the past decade, deep learning has exhibited remarkable empirical success across a wide range of engineering and scientific applications [24], such as computer vision [14, 43], natural language processing [48, 49], and health care [11], to name a few.
-As argued by Bengio et al.
-[2], Ma et al.
-[31], one major factor contributing to the success of deep learning is the ability of deep networks to perform powerful nonlinear feature learning by converting the data distribution to a compact and structured representation.
-This representation greatly facilitates various downstream tasks, including classification [10], segmentation [21], and generation [42].
-Based on the theory of data compression and optimal coding [30], Chan et al.
-[4], Yu et al.
-[56] proposed a principled and unified framework for deep learning to learn a compact and structured representation.
+In the past decade, deep learning has exhibited remarkable empirical success across a wide range of engineering and scientific applications [lecun2015deep], such as computer vision [he2016deep, simonyan2014very], natural language processing [sutskever2014sequence, vaswani2017attention], and health care [esteva2019guide], to name a few.
+As argued by bengio2013representation, ma2022principles, one major factor contributing to the success of deep learning is the ability of deep networks to perform powerful nonlinear feature learning by converting the data distribution to a compact and structured representation.
+This representation greatly facilitates various downstream tasks, including classification [dosovitskiy2020image], segmentation [kirillov2023segment], and generation [saharia2022photorealistic].
+Based on the theory of data compression and optimal coding [ma2007segmentation], chan2022redunet, yu2020learning proposed a principled and unified framework for deep learning to learn a compact and structured representation.
 Specifically, they proposed to maximize the difference between the coding rate of all features and the sum of coding rates of features in each class, which is referred to as maximal coding rate reduction (MCR2).
-This problem is presented in Problem (4) and visualized in Figure 1(a).
-Here, the coding rate measures the “compactness" of the features, which is interpreted as the volume of a particular set spanned by the learned features: a lower coding rate implies a more compact feature set [^1] Please refer to Chan et al.
-[4, Section 2.1] for more details on measuring compactness of feature sets via coding rates..
+This problem is presented in Problem (4) and visualized in Figure˜1(a).
+Here, the coding rate measures the “compactness" of the features, which is interpreted as the volume of a particular set spanned by the learned features: a lower coding rate implies a more compact feature set [^1] Please refer to chan2022redunet for more details on measuring compactness of feature sets via coding rates..
 Consequently, the MCR2 objective aims to maximize the volume of the set of all features while minimizing the volumes of the sets of features from each class.
-Motivated by the structural similarities between deep networks and unrolled optimization schemes for sparse coding [13, 34], Chan et al.
-[4] constructed a new deep network based on an iterative gradient descent scheme to maximize the MCR2 objective.
+Motivated by the structural similarities between deep networks and unrolled optimization schemes for sparse coding [gregor2010learning, monga2021algorithm], chan2022redunet constructed a new deep network based on an iterative gradient descent scheme to maximize the MCR2 objective.
 [^2] When performing maximization, we actually mean that we use gradient ascent.
@@
 Notably, each component of this deep network has precise optimization and geometric interpretations.
-Moreover, it has achieved strong empirical performance on various vision and language tasks [6, 57].
+Moreover, it has achieved strong empirical performance on various vision and language tasks [chu2023image, yu2023sparse].
 (a) Visualization of feature learning via MCR2.
@@
 The objective of MCR2 is to learn a feature mapping $f_{\bm{\Theta}}(\bm{x})\in\mathbb{R}^{d}$ such that $\bm{z}_{i}=f_{\bm{\Theta}}(\bm{x}_{i})$ for all $i$ are low-dimensional, discriminative, and diverse.
-(b) According to Theorem 1 and Theorem 2, the regularized MCR2 problem has a benign optimization landscape: each critical point is either a local maximizer or a strict saddle point.
+(b) According to Theorem˜1 and Theorem˜2, the regularized MCR2 problem has a benign optimization landscape: each critical point is either a local maximizer or a strict saddle point.
 Furthermore, each local maximizer, just like the global maximizer, corresponds to a feature representation that consists of a family of orthogonal subspaces, as illustrated in the middle.
@@
 Next, we show that the regularized MCR2 function possesses a benign global optimization landscape, despite its complicated structures.
-More precisely, each critical point is either a local maximizer or strict saddle point of the regularized MCR2 problem; see Figure 1(b).
+More precisely, each critical point is either a local maximizer or strict saddle point of the regularized MCR2 problem; see Figure˜1(b).
 Consequently, any gradient-based optimization, such as (stochastic) gradient descent, with random initialization can escape saddle points and at least converge to a local maximizer efficiently.
@@
 Namely, our theoretical characterizations of local and global optimality offer a compelling explanation for the empirical observations that both deep networks constructed via gradient descent applied to the MCR2 objective and over-parameterized deep networks trained by optimizing the MCR2 objective learn low-dimensional, discriminative, and diverse representations.
-These results align with the motivations of Chan et al.
-[4], Yu et al.
-[56] for employing the MCR2 principle for deep learning, and elucidate the outstanding performance of MCR2-based neural networks across a wide range of vision and language tasks [6, 59].
+These results align with the motivations of chan2022redunet, yu2020learning for employing the MCR2 principle for deep learning, and elucidate the outstanding performance of MCR2-based neural networks across a wide range of vision and language tasks [chu2023image, yu2023emergence].
 Moreover, our results underscore the potential of MCR2-based approaches to serve as a cornerstone for future advancements in deep learning, offering a principled approach to pursuing structured and compact representations in practical applications.
@@
 Low-dimensional structures in deep representation learning.
-In the literature, it has long been believed that the role of deep networks is to learn certain (nonlinear) low-dimensional and informative representations of the data [16, 31].
-For example, Papyan et al.
-[38] showed that the features learned by cross-entropy (CE) loss exhibit a neural collapse phenomenon during the terminal phase of training, where the features from the same class are mapped to a vector while the features from different classes are maximally linearly separable.
-Ansuini et al.
-[1], Recanatesi et al.
-[40] demonstrated that the dimension of the intermediate features first rapidly increases and then decreases from shallow to deep layers.
-Masarczyk et al.
-[32] concluded that the deep layers of neural networks progressively compress within-class features to learn low-dimensional features.
-Notably, Wang et al.
-[53] proposed a theoretical framework to analyze hierarchical feature learning for learning low-dimensional representations.
+In the literature, it has long been believed that the role of deep networks is to learn certain (nonlinear) low-dimensional and informative representations of the data [hinton2006reducing, ma2022principles].
+For example, papyan2020prevalence showed that the features learned by cross-entropy (CE) loss exhibit a neural collapse phenomenon during the terminal phase of training, where the features from the same class are mapped to a vector while the features from different classes are maximally linearly separable.
+ansuini2019intrinsic, recanatesi2019dimensionality demonstrated that the dimension of the intermediate features first rapidly increases and then decreases from shallow to deep layers.
+masarczyk2023tunnel concluded that the deep layers of neural networks progressively compress within-class features to learn low-dimensional features.
+Notably, wang2023understanding proposed a theoretical framework to analyze hierarchical feature learning for learning low-dimensional representations.
 They showed that each layer of deep linear networks progressively compresses within-class features and discriminates between-class features in classification problems.
 The MCR2-based approach to deep learning.
-The MCR2-based approach to deep learning for seeking structured and compact representations was first proposed by Yu et al.
-[56].
+The MCR2-based approach to deep learning for seeking structured and compact representations was first proposed by yu2020learning.
 Notably, they provided a global optimality analysis of the MCR2 problem (4) with additional rank constraints on the feature matrix of each class.
-Chan et al.
-[4] designed a new multi-layer deep network architecture, named ReduNet, based on an iterative gradient descent scheme for maximizing the MCR2 objective.
-To learn self-consistent representations, Dai et al.
-[8] extended this approach to the closed-loop transcription (CTRL) framework, which is formulated as a max-min game to optimize a modified MCR2 objective.
-This game was shown to have global equilibria corresponding to compact and structured representations [37].
-Recently, Yu et al.
-[58] showed that a transformer-like architecture named CRATE, which obtains strong empirical performance [6, 57, 59], can be naturally derived through an iterative optimization scheme for maximizing the sparse rate reduction objective, which is an adaptation to sequence data of the MCR2 objective studied in this work.
+chan2022redunet designed a new multi-layer deep network architecture, named ReduNet, based on an iterative gradient descent scheme for maximizing the MCR2 objective.
+To learn self-consistent representations, dai2022ctrl extended this approach to the closed-loop transcription (CTRL) framework, which is formulated as a max-min game to optimize a modified MCR2 objective.
+This game was shown to have global equilibria corresponding to compact and structured representations [pai2022pursuit].
+Recently, yu2023white showed that a transformer-like architecture named CRATE, which obtains strong empirical performance [chu2023image, yu2023sparse, yu2023emergence], can be naturally derived through an iterative optimization scheme for maximizing the sparse rate reduction objective, which is an adaptation to sequence data of the MCR2 objective studied in this work.
 Notation.
@@
 2 Problem Setup
-In this section, we first review the basic concepts of MCR2 for deep representation learning in Section 2.1, and then introduce our studied problem in Section 2.2.
+In this section, we first review the basic concepts of MCR2 for deep representation learning in Section˜2.1, and then introduce our studied problem in Section˜2.2.
 2.1 An Overview of MCR2
 In deep representation learning, given data $\{\bm{x}_{i}\}_{i=1}^{m}\subseteq\mathbb{R}^{n}$ from multiple (say $K$) classes, the goal is to learn neural-network representations of these samples that facilitate downstream tasks.
-Recent empirical studies have shown that good features can be learned for tasks such as classification or autoencoding by using heuristics to promote either the contraction of samples in the same class [41] or the contrast of samples between different classes [15, 36] during the training of neural networks.
-Notably, Chan et al.
-[4], Yu et al.
-[56] unified and formalized these practices and demonstrated that the MCR2 objective is an effective objective to learn within-class compressible and between-class discriminative representations of the data.
+Recent empirical studies have shown that good features can be learned for tasks such as classification or autoencoding by using heuristics to promote either the contraction of samples in the same class [rifai2011contractive] or the contrast of samples between different classes [he2020momentum, oord2018representation] during the training of neural networks.
+Notably, chan2022redunet, yu2020learning unified and formalized these practices and demonstrated that the MCR2 objective is an effective objective to learn within-class compressible and between-class discriminative representations of the data.
 The formulation of MCR2.
@@
 $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\mathbb{R}^{d\times m}.$ (1)
-On one hand, to make features between different classes discriminative or contrastive, one can maximize the lossy coding rate of all features in $\bm{Z}$, as argued in [4, 56], as follows:
+On one hand, to make features between different classes discriminative or contrastive, one can maximize the lossy coding rate of all features in $\bm{Z}$, as argued in [chan2022redunet, yu2020learning], as follows:
 $R(\bm{Z}):=\frac{1}{2}\log\det\left(\bm{I}+\frac{d}{m\epsilon^{2}}\bm{Z}\bm{Z}^{T}\right),$ (2)
 where $\epsilon>0$ is a prescribed quantization error.
-[^4] Here, $R(\bm{Z})$ is also known as the rate-distortion function in information theory [7], which represents the average number of binary bits needed to encode the data $\bm{Z}$.
+[^4] Here, $R(\bm{Z})$ is also known as the rate-distortion function in information theory [cover1999elements], which represents the average number of binary bits needed to encode the data $\bm{Z}$.
 On the other hand, to make features from the same class compressible or contractive, one can minimize the average lossy coding rate of features in the $k$-th class as follows:
@@
 Consequently, a good representation tends to maximize the difference between the coding rate for the whole and that for each class as follows:
-$\max_{\bm{Z}\in\mathbb{R}^{d\times m}}\ R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)\quad\mathrm{s.t.}\ \|\bm{Z}_{k}\|_{F}^{2}=m_{k},\quad\forall k\in[K].$ (4)
-This is referred to as the principle of maximal coding rate reduction in [4, 56].
+$\max_{\bm{Z}\in\mathbb{R}^{d\times m}}\ R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)\qquad\mathrm{s.t.}\ \|\bm{Z}_{k}\|_{F}^{2}=m_{k},\qquad\forall k\in[K].$ (4)
+This is referred to as the principle of maximal coding rate reduction in [chan2022redunet, yu2020learning].
 It is worth mentioning that this principle can be extended to self-supervised and even unsupervised learning settings, where we learn the label vectors $\{\bm{\pi}^{k}\}_{k=1}^{K}$ during training.
@@
 Therefore, we consider the Lagrangian formulation of (4).
-This can be viewed as a tight relaxation or even an equivalent problem of (4) whose optimal solutions agree under specific settings of the regularization parameter; see Proposition 1.
+This can be viewed as a tight relaxation or even an equivalent problem of (4) whose optimal solutions agree under specific settings of the regularization parameter; see ˜1.
 Specifically, the formulation we study, referred to henceforth as the regularized MCR2 problem, is as follows:
@@
 Applications of our formulation to deep representation learning via unrolled optimization.
-The first approach, as argued by Chan et al.
-[4], Yu et al.
-[57], is to construct a new deep network architecture, i.e., ReduNet [4] or CRATE [57], based on an iterative gradient descent scheme to optimize the MCR2-type objective.
+The first approach, as argued by chan2022redunet, yu2023sparse, is to construct a new deep network architecture, i.e., ReduNet [chan2022redunet] or CRATE [yu2023sparse], based on an iterative gradient descent scheme to optimize the MCR2-type objective.
 In this approach, each layer of the constructed network approximates a gradient descent step to optimize the MCR2-type objective given the input representation.
@@
 Applications of our formulation to deep representation learning with standard neural networks.
-In the second approach, one parameterizes the feature mapping $f_{\bm{\Theta}}(\cdot)$ via standard deep neural networks such as a multi-layer perceptron or ResNet [14], and treats the MCR2-type objective like other loss functions applied to outputs of a neural network, such as mean-squared error or cross-entropy loss.
+In the second approach, one parameterizes the feature mapping $f_{\bm{\Theta}}(\cdot)$ via standard deep neural networks such as a multi-layer perceptron or ResNet [he2016deep], and treats the MCR2-type objective like other loss functions applied to outputs of a neural network, such as mean-squared error or cross-entropy loss.
 Studying Problem (5) from this perspective would require us to optimize over $\bm{\Theta}$ instead of over $\bm{Z}$.
 This new optimization problem would be extraordinarily difficult to analyze, because modern neural networks have nonlinear interactions across many layers, so the parameters $\bm{\Theta}$ would affect the final representation $\bm{Z}$ in a complex way.
-Fortunately, since modern neural networks are often highly over-parameterized, they can interpolate or approximate any continuous function in the feature space [29], so we may omit these constraints by assuming the unconstrained feature model, where $\bm{z}_{i}$ for all $i\in[N]$ are treated as free optimization variables [33, 55, 62, 51].
+Fortunately, since modern neural networks are often highly over-parameterized, they can interpolate or approximate any continuous function in the feature space [lu2017expressive], so we may omit these constraints by assuming the unconstrained feature model, where $\bm{z}_{i}$ for all $i\in[N]$ are treated as free optimization variables [mixon2020neural, yaras2022neural, zhu2021geometric, wang2022linear].
 Consequently, studying the optimization properties of Problem (5) provides valuable insights into the structures of learned representations and the efficiency of training deep networks using MCR2-type objectives.
@@
 3 Main Results
-In this section, we first characterize the local and global optimal solutions of Problem (5) in Section 3.1, and then analyze the global landscape of the objective function in Section 3.2.
+In this section, we first characterize the local and global optimal solutions of Problem (5) in Section˜3.1, and then analyze the global landscape of the objective function in Section˜3.2.
 3.1 Characterization of Local and Global Optimality
@@
 Characterization of global maximizers) $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]$ is a global maximizer of Problem (5) if and only if (a) it satisfies the above all conditions and $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$, and (b) for all $k\neq l\in[K]$ satisfying $m_{k}<m_{l}$ and $r_{l}>0$, we have $r_{k}=\min\{m_{k},d\}$.
-We defer the proof to Section 4.1 and Section D.1.
+We defer the proof to Section˜4.1 and Section˜D.1.
 In this theorem, we explicitly characterize the local and global optima of Problem (5).
@@
 Our above discussion explains the merits of achieving both local and global optima.
-At each maximizer, the representations are within-class compressible and between-class discriminative (Theorem 1 (i)).
-Moreover, global maximizers further satisfy that the representations are all maximally diverse (Theorem 1 (ii)(a)).
-If all classes were balanced, i.e., $m_{1}=\cdots=m_{K}$, then Theorem 1 (ii)(b) would not apply, and these properties would be all that Theorem 1 asserts.
+At each maximizer, the representations are within-class compressible and between-class discriminative (Theorem˜1 (i)).
+Moreover, global maximizers further satisfy that the representations are all maximally diverse (Theorem˜1 (ii)(a)).
+If all classes were balanced, i.e., $m_{1}=\cdots=m_{K}$, then Theorem˜1 (ii)(b) would not apply, and these properties would be all that Theorem˜1 asserts.
 In this case, global optima would clearly be desired over local optima.
-However, in the unbalanced case, the situation is more complex, because Theorem 1 (ii)(b) would apply.
+However, in the unbalanced case, the situation is more complex, because Theorem˜1 (ii)(b) would apply.
 It says that for global optima, the classes with the smallest numbers of samples would fill to the largest dimension possible, and the very largest classes could collapse to $\bm{0}$, an undesirable situation.
@@
 In particular, local optima with more equitable rank distributions (like bigger classes span more dimensions) which are still maximally diverse (i.e., ranks of each class sum to the dimension $d$) could be preferred in applications.
-As demonstrated in Section 5.1, these kinds of potentially useful local optima are realized in experiments, even with unbalanced classes.
+As demonstrated in Section˜5.1, these kinds of potentially useful local optima are realized in experiments, even with unbalanced classes.
 Relation between Problems (4) and (5).
-Based on the characterization of global optimality in Theorem 1, we show the following proposition that establishes the relationship between the constrained MCR2 problem (4) and the regularized MCR2 problem (5) in terms of their global solutions under an appropriate choice of the regularization parameter.
-The proof of this result can be found in Section D.2.
+Based on the characterization of global optimality in Theorem˜1, we show the following proposition that establishes the relationship between the constrained MCR2 problem (4) and the regularized MCR2 problem (5) in terms of their global solutions under an appropriate choice of the regularization parameter.
+The proof of this result can be found in Section˜D.2.
 Proposition 1.
@@
 3.2 Analysis of Global Optimization Landscape
-While we have characterized the local and global optimal solutions in Theorem 1, it remains unknown whether these solutions can be computed efficiently using GD to solve Problem (5), as GD may get stuck at a saddle point.
-Fortunately, Sun et al.
-[44], Lee et al.
-[25] showed that if a function is twice continuously differentiable and satisfies strict saddle property, i.e., each critical point is either a local minimizer or a strict saddle point [^5] We say that a critical point is a strict saddle point of Problem (5) if it has a direction with strictly positive curvature; see Definition 2.
+While we have characterized the local and global optimal solutions in Theorem˜1, it remains unknown whether these solutions can be computed efficiently using GD to solve Problem (5), as GD may get stuck at a saddle point.
+Fortunately, sun2015nonconvex, lee2016gradient showed that if a function is twice continuously differentiable and satisfies strict saddle property, i.e., each critical point is either a local minimizer or a strict saddle point [^5] We say that a critical point is a strict saddle point of Problem (5) if it has a direction with strictly positive curvature; see Definition 2.
 This includes classical saddle points with strictly positive curvature as well as local minimizers., GD converges to its local minimizer almost surely with random initialization.
@@
 Given a coding precision $\epsilon>0$, if the regularization parameter satisfies (6), it holds that any critical point $\bm{Z}$ of Problem (5) that is not a local maximizer is a strict saddle point.
-We defer the proof to Section 4.2 and Section D.3. Here, we make some remarks on this theorem and also on the consequences of the results derived so far.
+We defer the proof to Section˜4.2 and Section˜D.3. Here, we make some remarks on this theorem and also on the consequences of the results derived so far.
 Differences from existing results on the MCR2 problem.
-Chan et al.
-[4], Yu et al.
-[56] have characterized the global optimality of Problem (4) with Frobenius norm constraints on each $\bm{Z}_{k}$ in the UFM.
+chan2022redunet, yu2020learning have characterized the global optimality of Problem (4) with Frobenius norm constraints on each $\bm{Z}_{k}$ in the UFM.
 However, their analysis requires an additional rank constraint on each $\bm{Z}_{k}$ and only characterizes globally optimal representations.
@@
 Furthermore, we demonstrate that the regularized MCR2 objective (5) is a strict saddle function.
-To the best of our knowledge, Theorems 1 and 2 constitute the first analysis of local optima and optimization landscapes for MCR2 objectives.
-According to Daneshmand et al.
-[9], Lee et al.
-[25], Xu et al.
-[54], Theorems 1 and 2 imply that low-dimensional and discriminative representations can be efficiently found by (stochastic) GD on Problem (5) from a random initialization.
+To the best of our knowledge, Theorems˜1 and 2 constitute the first analysis of local optima and optimization landscapes for MCR2 objectives.
+According to daneshmand2018escaping, lee2016gradient, xu2018first, Theorems˜1 and 2 imply that low-dimensional and discriminative representations can be efficiently found by (stochastic) GD on Problem (5) from a random initialization.
 Comparison to existing landscape analyses in non-convex optimization.
 In recent years, there has been a growing body of literature exploring optimization landscapes of non-convex problems in machine learning and deep learning.
-These include low-rank matrix factorization [12, 46, 5, 60], community detection [50, 52], dictionary learning [45, 39], and deep neural networks [47, 55, 61, 62, 18, 27].
+These include low-rank matrix factorization [ge2017no, sun2018geometric, chi2019nonconvex, zhang2020symmetry], community detection [wang2021optimal, wang2022non], dictionary learning [sun2017complete_a, qu2020geometric], and deep neural networks [sun2020global, yaras2022neural, zhou2022optimization, zhu2021geometric, jiang2024generalized, li2024neural].
 The existing analyses in the literature cannot be applied to the MCR2 problem due to its special structure, which involves the log-determinant of all features minus the sum of the log-determinant of features in each class.
@@
 4 Proofs of Main Results
-In this section, we sketch the proofs of our main theorems in Section 3. The complete proofs can be found in Sections B and C of the appendix. For ease of exposition, let
-$\alpha:=\frac{d}{m\epsilon^{2}},\quad\alpha_{k}:=\frac{d}{m_{k}\epsilon^{2}},\quad\forall k\in[K].$ (11)
+In this section, we sketch the proofs of our main theorems in Section˜3. The complete proofs can be found in Sections B and C of the appendix. For ease of exposition, let
+$\alpha:=\frac{d}{m\epsilon^{2}},\qquad\alpha_{k}:=\frac{d}{m_{k}\epsilon^{2}},\qquad\forall k\in[K].$ (11)
 4.1 Analysis of Optimality Conditions
@@
 In particular, this upper bound is tight when the blocks $\{\bm{Z}_{k}\}_{k=1}^{K}$ are orthogonal to each other.
-This result is a direct consequence of [4, Lemma 10].
+This result is a direct consequence of [chan2022redunet, Lemma 10].
 Lemma 1.
@@
 Proposition 2.
-Consider the setting of Theorem 1. It holds that $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\cal Z$ if and only if each $\bm{Z}_{k}$ admits the following singular value decomposition
+Consider the setting of Theorem˜1. It holds that $\bm{Z}=\left[\bm{Z}_{1},\dots,\bm{Z}_{K}\right]\in\cal Z$ if and only if each $\bm{Z}_{k}$ admits the following singular value decomposition
 $\bm{Z}_{k}=\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}\bm{V}_{k}^{T},\ \tilde{\bm{\Sigma}}_{k}=\mathrm{diag}\left(\sigma_{k,1},\dots,\sigma_{k,r_{k}}\right),$ (14)
@@
 where $\eta_{k}=(\alpha_{k}-\alpha)-\lambda\left(m/m_{k}+1\right)$ and
-$\overline{\sigma}_{k}=\left(\frac{\eta_{k}+\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2},\quad\underline{\sigma}_{k}=\left(\frac{\eta_{k}-\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2}.$ (16)
+$\overline{\sigma}_{k}=\left(\frac{\eta_{k}+\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2},\qquad\underline{\sigma}_{k}=\left(\frac{\eta_{k}-\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}\right)^{1/2}.$ (16)
 This proposition shows that each critical point that is between-class discriminative (i.e., $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$) exhibits a specific structure: the singular values of $\bm{Z}_{k}$ can only take on two possible values, $\overline{\sigma}_{k}$ and $\underline{\sigma}_{k}$.
@@
 $\mathcal{Z}^{c}:=\left\{\bm{Z}:\nabla F(\bm{Z})=0,\ \bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0},\ \exists k\neq l\right\}.$ (18)
-Moreover, according to Proposition 2, we further divide $\cal Z$ into two disjoint sets $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Here,
+Moreover, according to ˜2, we further divide $\cal Z$ into two disjoint sets $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Here,
 $\mathcal{Z}_{1}:=\mathcal{Z}\cap\left\{\bm{Z}:\sigma_{k,i}(\bm{Z}_{k})=\overline{\sigma}_{k},\forall i\in[r_{k}],k\in[K]\right\},\ \mathcal{Z}_{2}:=\mathcal{Z}\setminus\mathcal{Z}_{1},$ (19)
@@
 Proposition 3.
-Consider the setting of Theorem 2. Suppose that $\bm{Z}\in\cal Z$. Then, the following statements hold:
+Consider the setting of Theorem˜2. Suppose that $\bm{Z}\in\cal Z$. Then, the following statements hold:
 (i) If $\bm{Z}_{k}$ takes the form of (14) with $\sigma_{k,i}=\overline{\sigma}_{k}$ for all $i\in[r_{k}]$ and all $k\in[K]$, i.e., $\bm{Z}\in\mathcal{Z}_{1}$, then $\bm{Z}$ is a local maximizer.
@@
 Next, we proceed to the second step to show that any point belonging to $\mathcal{Z}^{c}$ is a strict saddle point.
-It suffices to find a direction $\bm{D}\in\mathbb{R}^{d\times m}$ such that $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0$ for each $\bm{Z}\in\mathcal{Z}^{c}$ according to Definition 2.
+It suffices to find a direction $\bm{D}\in\mathbb{R}^{d\times m}$ such that $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]>0$ for each $\bm{Z}\in\mathcal{Z}^{c}$ according to Definition˜2.
 Proposition 4.
-Consider the setting of Theorem 2.
+Consider the setting of Theorem˜2.
 If $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point and there exists $1\leq k\neq l\leq K$ such that $\bm{Z}_{k}^{T}\bm{Z}_{l}\neq\bm{0}$, i.e., $\bm{Z}\in\mathcal{Z}^{c}$, then $\bm{Z}$ is a strict saddle point.
-With the above preparations that characterize all the critical points, we can prove Theorem 1 and Theorem 2. We refer the reader to Appendix D for the detailed proof.
+With the above preparations that characterize all the critical points, we can prove Theorem˜1 and Theorem˜2. We refer the reader to Appendix˜D for the detailed proof.
 5 Experimental Results
@@
 The number of nonzero singular values in each subspace is $24,23,27,26$, respectively.
-In this section, we first conduct numerical experiments on synthetic data in Section 5.1 to validate our theoretical results, and then on real-world data sets using deep neural networks in Section 5.2 to further support our theory.
+In this section, we first conduct numerical experiments on synthetic data in Section˜5.1 to validate our theoretical results, and then on real-world data sets using deep neural networks in Section˜5.2 to further support our theory.
 All codes are implemented in Python mainly using NumPy and PyTorch.
 All of our experiments are executed on a computing server equipped with NVIDIA A40 GPUs.
-Due to space limitations, we defer some implementation details and additional experimental results to Appendix E.
+Due to space limitations, we defer some implementation details and additional experimental results to Appendix˜E.
 5.1 Validation of Theory for Solving Problem (5)
@@
 We visualize the optimization dynamics and structures of the solutions returned by GD to verify and validate Theorems 1 and 2.
-Verification of Theorem 1.
+Verification of Theorem˜1.
 In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=30,m_{2}=70,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$.
 Then, one can verify that $\lambda$ satisfies (6).
-For the solution $\bm{Z}$ returned by GD, we first plot the heatmap of the cosine similarity between pairwise columns of $\bm{Z}$ in Figure 2(a).
+For the solution $\bm{Z}$ returned by GD, we first plot the heatmap of the cosine similarity between pairwise columns of $\bm{Z}$ in Figure˜2(a).
 We observe that the features from different classes are orthogonal to each other, while the features from the same class are correlated.
-Next, we compute the singular values of $\bm{Z}_{k}$ via singular value decomposition (SVD) and plot the singular values using blue dots for each $k\in[K]$ in Figure 2(b).
-According to the closed-form solution (7) in Theorem 1, we also plot the theoretical bound of singular values in red in Figure 2(b).
+Next, we compute the singular values of $\bm{Z}_{k}$ via singular value decomposition (SVD) and plot the singular values using blue dots for each $k\in[K]$ in Figure˜2(b).
+According to the closed-form solution (7) in Theorem˜1, we also plot the theoretical bound of singular values in red in Figure˜2(b).
 One can observe that the number of singular values of each block is respectively $24,23,27,26$, summing up to $100$, and the red line perfectly matches the blue dots.
-These results all provide strong support for Theorem 1.
-Verification of Theorem 2.
+These results all provide strong support for Theorem˜1.
+Verification of Theorem˜2.
 In this experiment, we maintain the same setting as above, except that the number of samples in each class is equal.
 We first fix $m=200$ and vary $d\in\{40,80,120\}$, and then fix $d=50$ and vary $d\in\{100,200,400\}$ to run GD.
-We plot the distances between function values of the iterates to the optimal value, which is computed according to (7) in Theorem 1, against the iteration numbers in Figure 3.
+We plot the distances between function values of the iterates to the optimal value, which is computed according to (7) in Theorem˜1, against the iteration numbers in Figure˜3.
 We observe that GD with random initialization converges to an optimal solution at a linear rate.
-This indicates that the MCR2 has a benign global landscape, which supports Theorem 2.
+This indicates that the MCR2 has a benign global landscape, which supports Theorem˜2.
 (a) Convergence for $m=200,d\in\{40,80,120\}$
@@
 Figure 3: Convergence performance of GD for solving the regularized MCR2 problem.
-Here, the $x$-axis is number of iterations (also denoted by $t$), and $y$-axis is the function value gap $F^{t}-F^{*}$, where $F^{t}=F(\bm{Z}^{t})$ denotes the function value at the $t$-th iterate $\bm{Z}^{t}$ generated by GD, and $F^{*}$ is the optimal value of Problem (5) computed according to (7) in Theorem 1.
+Here, the $x$-axis is number of iterations (also denoted by $t$), and $y$-axis is the function value gap $F^{t}-F^{*}$, where $F^{t}=F(\bm{Z}^{t})$ denotes the function value at the $t$-th iterate $\bm{Z}^{t}$ generated by GD, and $F^{*}$ is the optimal value of Problem (5) computed according to (7) in Theorem˜1.
 5.2 Training Deep Networks Using Regularized MCR2
-In this subsection, we conduct numerical experiments on the image datasets MNIST [23] and CIFAR-10 [22] to provide evidence that our theory also applies to deep networks.
+In this subsection, we conduct numerical experiments on the image datasets MNIST [lecun1998gradient] and CIFAR-10 [krizhevsky2009learning] to provide evidence that our theory also applies to deep networks.
 More specifically, we employ a multi-layer perceptron network with ReLU activation as the feature mapping $\bm{z}=f_{\bm{\Theta}}(\bm{x})$ with output dimension 32 for MNIST and 128 for CIFAR-10.
-Then, we train the network parameters $\bm{\Theta}$ via Adam [20] by optimizing Problem (5).
+Then, we train the network parameters $\bm{\Theta}$ via Adam [kingma2014adam] by optimizing Problem (5).
 Experimental setting and results.
@@
 We set $\lambda=0.001$ and $\epsilon=0.5$.
-For different subsets with corresponding values of $(m,K)$, we run experiments and report the function value $\hat{F}$ obtained by training deep networks and the optimal value $F^{*}$ computed using the closed-form solution in Theorem 1 in Table 1.
-To verify the discriminative nature of the features obtained by training deep networks across different classes, we measure the discrimination between features belonging to different classes by computing the cosine of the principal angle [3] between the class subspaces: $s=\max\left\{\|\bm{U}_{k}^{T}\bm{U}_{l}\|:k\neq l\in[K]\right\}\in[0,1]$, where the columns of $\bm{U}_{k}\in\mathbb{R}^{d\times r_{k}}$ are the right singular vectors corresponding to the top $r_{k}$ singular values of $\bm{Z}_{k}$ defined in (14) and $r_{k}$ is its rank [^6] We estimate the rank of a matrix by rounding its “stable rank” [17]: $r_{k}=\operatorname{round}(\|\bm{Z}_{k}\|_{F}^{2}/\|\bm{Z}_{k}\|^{2})$.
+For different subsets with corresponding values of $(m,K)$, we run experiments and report the function value $\hat{F}$ obtained by training deep networks and the optimal value $F^{*}$ computed using the closed-form solution in Theorem˜1 in Table˜1.
+To verify the discriminative nature of the features obtained by training deep networks across different classes, we measure the discrimination between features belonging to different classes by computing the cosine of the principal angle [bjorck1973numerical] between the class subspaces: $s=\max\left\{\|\bm{U}_{k}^{T}\bm{U}_{l}\|:k\neq l\in[K]\right\}\in[0,1]$, where the columns of $\bm{U}_{k}\in\mathbb{R}^{d\times r_{k}}$ are the right singular vectors corresponding to the top $r_{k}$ singular values of $\bm{Z}_{k}$ defined in (14) and $r_{k}$ is its rank [^6] We estimate the rank of a matrix by rounding its “stable rank” [horn2012matrix]: $r_{k}=\operatorname{round}(\|\bm{Z}_{k}\|_{F}^{2}/\|\bm{Z}_{k}\|^{2})$.
 for each $k\in[K]$.
 In particular, when $s$ is smaller, the spaces spanned by each pair $\bm{Z}_{k}$ and $\bm{Z}_{l}$ for $k\neq l$ are closer to being orthogonal to each other.
-Then, we record the value $s$ in Table 1 in different settings.
-Moreover, we visualize the pairwise cosine similarities between learned features on MNIST and CIFAR-10 when $(m,K)=(1500,6)$ and $(2500,10)$ in Figure 4.
+Then, we record the value $s$ in Table˜1 in different settings.
+Moreover, we visualize the pairwise cosine similarities between learned features on MNIST and CIFAR-10 when $(m,K)=(1500,6)$ and $(2500,10)$ in Figure˜4.
 Table 1: Function value $\hat{F}$ obtained by training deep networks, the optimal value $F^{*}$ computed by our theory on subsets of MNIST or CIFAR-10, and discrimination metric $s$ of features.
@@
 (2500, 10) 228.48 228.49 0.171
-We observe from Table 1 that the function value returned by training deep networks is extremely close to the global optimal value of Problem (5) and from the value $s$ and Figure 4 that the features from different classes are nearly orthogonal to each other.
-These observations, together with Theorems 1 and 2, indicate that Problem (5) retains its optimization properties even when $\bm{Z}$ is parameterized by a neural network.
+We observe from Table˜1 that the function value returned by training deep networks is extremely close to the global optimal value of Problem (5) and from the value $s$ and Figure˜4 that the features from different classes are nearly orthogonal to each other.
+These observations, together with Theorems˜1 and 2, indicate that Problem (5) retains its optimization properties even when $\bm{Z}$ is parameterized by a neural network.
 Our theoretical analysis of Problem (5) thus illustrates a qualitative picture of training deep networks with the regularized MCR2 objective.
@@
 For future work, it is natural to extend our analysis to Problem (4) with deep network parameterizations.
-It is also interesting to study the sparse MCR2 objective, which has led to high-performance transformer-like architectures [57, 58].
+It is also interesting to study the sparse MCR2 objective, which has led to high-performance transformer-like architectures [yu2023sparse, yu2023white].
 Acknowledgements
@@
 Supplementary Material
-The organization of the supplementary material is as follows: In Appendix A, we introduce preliminary setups and auxiliary results for studying the MCR2 problem.
-Then, we prove the technical results concerning the global optimality of Problem (5) in Appendix B and the optimization landscape of Problem (5) in Appendix C, respectively.
-In Appendix D, we prove the main theorems in Theorem 1 and Theorem 2.
-Finally, we provide more experimental setups and results in Appendix E.
+The organization of the supplementary material is as follows: In Appendix˜A, we introduce preliminary setups and auxiliary results for studying the MCR2 problem.
+Then, we prove the technical results concerning the global optimality of Problem (5) in Appendix˜B and the optimization landscape of Problem (5) in Appendix˜C, respectively.
+In Appendix˜D, we prove the main theorems in Theorem˜1 and Theorem˜2.
+Finally, we provide more experimental setups and results in Appendix˜E.
 Besides the notions introduced earlier, we shall use $\mathrm{BlkDiag}(\bm{X}_{1},\dots,\bm{X}_{K})$ to denote the block diagonal matrix whose diagonal blocks are $\bm{X}_{1},\dots,\bm{X}_{K}$.
@@
 $\nabla^{2}R(\bm{Z})[\bm{D},\bm{D}]$ $=\alpha\langle\bm{X}^{-1},\bm{D}\bm{D}^{T}\rangle-\frac{\alpha^{2}}{2}\mathrm{Tr}\left(\bm{X}^{-1}(\bm{Z}\bm{D}^{T}+\bm{D}\bm{Z}^{T})\bm{X}^{-1}(\bm{Z}\bm{D}^{T}+\bm{D}\bm{Z}^{T})\right),$ (23)
-where $\bm{X}:=\bm{I}_{n}+\alpha\bm{Z}\bm{Z}^{T}$ and $\alpha$ is defined in (11).
+where $\bm{X}:=\bm{I}_{d}+\alpha\bm{Z}\bm{Z}^{T}$ and $\alpha$ is defined in (11).
 Note that we can compute the gradient and Hessian of $R_{c}(\cdot)$ in (20) using the same approach.
@@
 where $\alpha$ and $\alpha_{k}$ are defined in (11).
-According to Jin et al.
-[19], Lee et al.
-[26], we define the strict saddle point, i.e., a critical point that has a direction with strictly positive curvature [^7] Note that Problem (5) is not a minimization problem but a maximization problem., of Problem (5) as follows:
+According to jin2017escape, lee2019first, we define the strict saddle point, i.e., a critical point that has a direction with strictly positive curvature [^7] Note that Problem (5) is not a minimization problem but a maximization problem., of Problem (5) as follows:
 Definition 2.
@@
 $\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}]=\nabla^{2}R(\bm{Z})[\bm{D},\bm{D}]-\sum_{k=1}^{K}\frac{m_{k}}{m}\nabla^{2}R_{c}(\bm{Z}_{k})[\bm{D}_{k},\bm{D}_{k}]-\lambda\|\bm{D}\|_{F}^{2}.$ (25)
-Remark that for the MCR2 problem, strict saddle points include classical saddle points with strictly positive curvature as well as local minimizers.
+Remark that for the MCR2 problem, strict saddle points include saddle points with strictly positive curvature as well as local maximizers.
 A.2 Auxiliary Results
@@
 We next present the commutative property for the log-determinant function and the upper bound for the coding rate function.
-We refer the reader to [4, Lemma 8 & Lemma 10] for the detailed proofs.
+We refer the reader to [chan2022redunet, Lemma 8 & Lemma 10] for the detailed proofs.
 Here, let $\bm{Z}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decompositon of $\bm{Z}\in\mathbb{R}^{d\times m}$, where $r=\mathrm{rank}(\bm{Z})\leq\min\{m,d\}$, $\bm{U}\in\mathcal{O}^{d\times r}$, $\bm{\Sigma}\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $\bm{V}\in\mathcal{O}^{m\times r}$.
@@
 For any $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$, where $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for each $k\in[K]$, we have
-$F(\bm{Z}\bm{O})=F(\bm{Z}),\quad\nabla F(\bm{Z}\bm{O})=\nabla F(\bm{Z})\bm{O},\quad\nabla^{2}F(\bm{Z}\bm{O})[\bm{D}\bm{O},\bm{D}\bm{O}]=\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}].$ (29)
+$F(\bm{Z}\bm{O})=F(\bm{Z}),\qquad\nabla F(\bm{Z}\bm{O})=\nabla F(\bm{Z})\bm{O},\qquad\nabla^{2}F(\bm{Z}\bm{O})[\bm{D}\bm{O},\bm{D}\bm{O}]=\nabla^{2}F(\bm{Z})[\bm{D},\bm{D}].$ (29)
+Proof of Lemma˜5.
+[…]
+Appendix B Proofs in Section˜4.1
+B.1 Proof of Lemma˜1
 Proof of Lemma 1.
-[…]
-Appendix B Proofs in Section 4.1
-B.1 Proof of Lemma 1
-Proof of Lemma 1.
-It follows from (28) in Lemma 4 that
+It follows from (28) in Lemma˜4 that
 $\log\det\left(\bm{I}_{d}+\alpha\bm{Z}\bm{Z}^{T}\right)\leq\sum_{k=1}^{K}\log\det\left(\bm{I}_{d}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right),$ (30)
 where the equality holds if and only if $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $1\leq k\neq l\leq K$. Substituting this into (5) directly yields (12). ∎
-B.2 Proof of Proposition 2
-Proof of Proposition 2.
+B.2 Proof of ˜2
+Proof of ˜2.
 Let $\bm{Z}\in\cal Z$ be arbitrary, where $\mathcal{Z}$ is defined in (13).
@@
 According to (6), one can verify that for each $k\in[K]$,
-$\eta_{k}>0,\quad\eta_{k}^{2}-\frac{4\alpha_{k}}{\alpha}\lambda^{2}\geq 0.$
+$\eta_{k}>0,\qquad\eta_{k}^{2}-\frac{4\alpha_{k}}{\alpha}\lambda^{2}\geq 0.$
 This yields that the above quadratic equation has positive roots as follows. For each $i\in[r_{k}]$ and $k\in[K]$, we have
 $\sigma_{k,i}^{2}=\frac{\eta_{k}\pm\sqrt{\eta_{k}^{2}-4\lambda^{2}m/m_{k}}}{2\lambda\alpha_{k}}.$ (37)
-Finally, using $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ and (33), we obtain $\bm{Q}_{k,1}^{T}\bm{Q}_{l,1}=\bm{0}$ for all $1\leq l\neq k\leq K$. These, together with (33), yields (14).
+Finally, using $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ and (33), we obtain $\bm{P}_{k,1}^{T}\bm{P}_{l,1}=\bm{0}$ for all $1\leq l\neq k\leq K$. These, together with (33), yields (14).
 Conversely, suppose that each block $\bm{Z}_{k}$ of $\bm{Z}$ satisfies $\bm{Z}_{k}=\bm{0}$ or takes the form (14) for some $\bm{U}_{k}\in\mathcal{O}^{d\times r_{k}}$ satisfying $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $l\neq k$, $\bm{V}_{k}\in\mathcal{O}^{m_{k}\times r_{k}}$ for all $k\in[K]$, and $\sigma_{k,i}>0$ satisfying (15).
@@
 where the last equality follows from (14) and (B.2). Therefore, we have $\nabla F(\bm{Z})=\bm{0}$ as desired. This, together with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$, implies $\bm{Z}\in\cal Z$. ∎
-Appendix C Proofs in Section 4.2
-C.1 Proof of Proposition 3
-Proof of Proposition 3.
-For each $\bm{Z}\in\cal Z$, it follows from Lemma 1 that
+Appendix C Proofs in Section˜4.2
+C.1 Proof of ˜3
+Proof of ˜3.
+For each $\bm{Z}\in\cal Z$, it follows from Lemma˜1 that
 $F(\bm{Z})=\sum_{k=1}^{K}f_{k}(\bm{Z}_{k}),$ (40)
@@
 Suppose to the contrary that $r_{k}>0$ for all $k\in[K]$.
-For each $\bm{Z}\in\cal Z$, using Lemma 1 with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$, (14), and (31), we have
+For each $\bm{Z}\in\cal Z$, using Lemma˜1 with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$, (14), and (31), we have
 $F(\bm{Z})$ $=\sum_{k=1}^{K}\left(\frac{1}{2}\log\det\left(\bm{I}_{n}+\alpha\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}^{2}\bm{U}_{k}^{T}\right)-\frac{m_{k}}{2m}\log\det\left(\bm{I}_{n}+\alpha_{k}\bm{U}_{k}\tilde{\bm{\Sigma}}_{k}^{2}\bm{U}_{k}^{T}\right)-\frac{\lambda}{2}\|\bm{Z}_{k}\|_{F}^{2}\right)$
@@
 This, together with (C.1) and the fact that $\bm{0}$ is a local maximizer of $f_{k}(\bm{Z}_{k})$, implies (i) and (ii).
-C.2 Proof of Proposition 4
-Proof of Proposition 4.
+C.2 Proof of ˜4
+Proof of ˜4.
 Note that $\bm{Z}\in\mathbb{R}^{d\times m}$ is a critical point that satisfies (24).
@@
 where $\bm{Q}_{i}\in\mathcal{O}^{d\times h_{i}}$ for all $i\in[p]$.
-According to Lemma 5, we can see that $\bm{Z}$ is a critical point with curvature if and only if $\bm{Z}\bm{O}$ is a critical point with the same curvature for each $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$ with $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for all $k\in[K]$.
+According to Lemma˜5, we can see that $\bm{Z}$ is a critical point with curvature if and only if $\bm{Z}\bm{O}$ is a critical point with the same curvature for each $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$ with $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for all $k\in[K]$.
 According to the SVD of $\bm{Z}_{k}$ in (33), we can take $\bm{O}_{k}=\bm{Q}_{k}$ for each $k\in[K]$.
 Therefore, it suffices to study $\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}$ for each $k\in[K]$.
-Substituting this into (24) in Definition 1 gives
+Substituting this into (24) in Definition˜1 gives
 $\alpha(\bm{I}+\alpha\bm{Z}\bm{Z}^{T})^{-1}\bm{P}_{k}\bm{\Sigma}_{k}-\alpha\bm{P}_{k}(\bm{I}+\alpha_{k}\bm{\Sigma}_{k}\bm{\Sigma}_{k}^{T})^{-1}\bm{\Sigma}_{k}-\lambda\bm{P}_{k}\bm{\Sigma}_{k}=\bm{0},\ \forall k\in[K].$
@@
 $\bm{Z}\bm{Z}^{T}=\sum_{k=1}^{K}\bm{Z}_{k}\bm{Z}_{k}^{T}=\sum_{k=1}^{K}\sum_{i=1}^{p}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}=\sum_{i=1}^{p}\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}.$
-This, together with (2), Lemma 4, and $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$, yields that
+This, together with (2), Lemma˜4, and $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$, yields that
 $R(\bm{Z})=\frac{1}{2}\sum_{i=1}^{p}\log\det\left(\bm{I}+\alpha\bm{Z}^{(i)}\bm{Z}^{(i)^{T}}\right).$ (47)
@@
 $=\alpha\|\bm{q}_{k}\|^{2}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}$
-$\quad-\alpha\alpha_{k}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)\frac{\beta_{i}^{2}\|\bm{q}_{k}\|^{2}}{\alpha_{k}\beta_{i}^{2}+1},$
+$\qquad-\alpha\alpha_{k}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)\frac{\beta_{i}^{2}\|\bm{q}_{k}\|^{2}}{\alpha_{k}\beta_{i}^{2}+1},$
 where $\bm{X}_{k}^{(i)}=\left(\bm{I}+\alpha_{k}\bm{Z}_{k}^{(i)}\bm{Z}_{k}^{(i)^{T}}\right)^{-1}$, the second equality follows from
@@
 Summing up the above equality for all $k\in[K]$ with $\alpha_{k}=K\alpha$ for all $k\in[K]$ yields
-$\quad\sum_{k=1}^{K}\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$
+$\qquad\sum_{k=1}^{K}\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$
 $=\alpha\left(1-\frac{\alpha_{k}\beta_{i}^{2}}{\alpha_{k}\beta_{i}^{2}+1}\right)\|\bm{q}\|^{2}-\frac{\alpha\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
-$\quad-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\frac{\alpha\alpha_{k}^{2}\beta_{i}^{2}}{\left(\alpha_{k}\beta_{i}^{2}+1\right)^{2}}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
+$\qquad-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\frac{\alpha\alpha_{k}^{2}\beta_{i}^{2}}{\left(\alpha_{k}\beta_{i}^{2}+1\right)^{2}}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
 $=\left(\frac{\alpha}{1+\alpha\lambda_{i}}-\lambda\right)\|\bm{q}\|^{2}-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right),$
@@
 where $\bm{Q}_{i}\in\mathcal{O}^{d\times h_{i}}$ for all $i\in[p]$.
-Appendix D Proofs in Section 3
-D.1 Proof of Theorem 1
+Appendix D Proofs in Section˜3
+D.1 Proof of Theorem˜1
 Proof of Theorem 1.
 (i) Suppose that each block of $\bm{Z}$ satisfying (7), (a), (b), and (c).
-It directly follows from (i) in Proposition 3 that $\bm{Z}$ is a local maximizer.
+It directly follows from (i) in ˜3 that $\bm{Z}$ is a local maximizer.
 Conversely, suppose that $\bm{Z}$ is a local maximizer.
-According to (17), (18), (19), (ii) in Proposition 3 and Proposition 4, if $\bm{Z}\in\mathcal{Z}^{c}\cup\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point.
+According to (17), (18), (19), (ii) in ˜3 and ˜4, if $\bm{Z}\in\mathcal{Z}^{c}\cup\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point.
 This, together with $\mathcal{X}=\mathcal{Z}^{c}\cup\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$ and the fact that $\bm{Z}$ is a local maximizer, implies that $\bm{Z}\in\mathcal{Z}_{1}$.
-Using this, (19), and Proposition 2 yields that $\bm{Z}$ satisfying (7), (a), (b), and (c).
+Using this, (19), and ˜2 yields that $\bm{Z}$ satisfying (7), (a), (b), and (c).
 (ii) According to (i) in Theorem 1, suppose that the $k$-th block of a local maximizer $\bm{Z}$ admits the decomposition in (7) satisfying (a), (b), and (c) for all $k\in[K]$.
-This, together with (C.1) in the proof of Proposition 3, yields that
+This, together with (C.1) in the proof of ˜3, yields that
 $F(\bm{Z})$ $=\frac{1}{2}\sum_{k=1}^{K}\sum_{i=1}^{r_{k}}\left(\log\left(1+\alpha\overline{\sigma}_{k}^{2}\right)-\frac{m_{k}}{m}\log\left(1+\alpha_{k}\overline{\sigma}_{k}^{2}\right)-\lambda\overline{\sigma}_{k}^{2}\right),$ (60)
@@
 ∎
-D.2 Proof of Proposition 1
-To prove Proposition 1, we first need to characterize the global optimal solution set of Problem (4).
+D.2 Proof of ˜1
+To prove ˜1, we first need to characterize the global optimal solution set of Problem (4).
 Proposition 1.
@@
 Proof.
-According to Lemma 4, we have
+According to Lemma˜4, we have
 $R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)$ $\leq\frac{1}{2}\sum_{k=1}^{K}\log\det\left(\bm{I}+\alpha\bm{Z}_{k}\bm{Z}_{k}^{T}\right)-\sum_{k=1}^{K}\frac{m_{k}}{2m}\log\det\left(\bm{I}+\alpha_{k}\bm{Z}_{k}\bm{Z}_{k}^{T}\right)$
@@
 Substituting this SVD into (63), together with $\|\bm{Z}_{k}\|_{F}^{2}=m_{k}$, yields that to maximize $H(\bm{Z})$, it suffices to study for each $k\in[K]$,
-$\max_{\sigma_{k,1},\dots,\sigma_{k,r_{k}}}\sum_{i=1}^{r_{k}}\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\sum_{i=1}^{r_{k}}\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\qquad\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}\sigma_{k,i}^{2}=m_{k}.$
+$\max_{\sigma_{k,1},\dots,\sigma_{k,r_{k}}}\sum_{i=1}^{r_{k}}\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\sum_{i=1}^{r_{k}}\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\hskip 18.49988pt\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}\sigma_{k,i}^{2}=m_{k}.$
 To simplify our development, let $x_{i}:=\sigma_{k,i}^{2}\geq 0$ for each $i\in[r_{k}]$. This, together with $m_{k}=m/K$, implies that it suffices to study
-$\max_{x_{1},\dots,x_{r_{k}}}g(\bm{x})$ $:=\sum_{i=1}^{r_{k}}\log\left(1+\alpha x_{i}\right)-\sum_{i=1}^{r_{k}}\frac{1}{K}\log\left(1+\alpha_{k}x_{i}\right)\qquad$
+$\max_{x_{1},\dots,x_{r_{k}}}g(\bm{x})$ $:=\sum_{i=1}^{r_{k}}\log\left(1+\alpha x_{i}\right)-\sum_{i=1}^{r_{k}}\frac{1}{K}\log\left(1+\alpha_{k}x_{i}\right)\hskip 18.49988pt$
 $\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}x_{i}=\frac{m}{K},\ \ x_{i}\geq 0,\ \forall i\in[r_{k}].$ (64)
-This, together with Lemma 6 and (11), yields that the optimal solution for each $k\in[K]$ is
+This, together with Lemma˜6 and (11), yields that the optimal solution for each $k\in[K]$ is
 $x_{i}^{*}=\frac{m}{r_{k}K},\ \forall i\in[r_{k}].$ (65)
@@
 Then, we have $r_{k}\leq m/K$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq m$.
-This, together with (65) and Lemma 7, implies that $r_{k}=m/K$, and thus $x_{i}^{*}=1$ for all $i\in[r_{k}]$ and $k\in[k]$.
+This, together with (65) and Lemma˜7, implies that $r_{k}=m/K$, and thus $x_{i}^{*}=1$ for all $i\in[r_{k}]$ and $k\in[k]$.
 (ii) Suppose that $m>d$. Then, we have $r_{k}\leq\min\{d,m/K\}$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq d$. To compute the optimal function value, we consider the following problem:
@@
 Then, we complete the proof.
-Based on the above proposition, we are ready to prove Proposition 1.
-Proof of Proposition 1.
-Let $\hat{\bm{Z}}=[\bm{Z}_{1},\dots,\bm{Z}_{k}]$ denote the optimal solution of Problem (4). According to Proposition 5, it suffices to study the following two cases.
+Based on the above proposition, we are ready to prove ˜1.
+Proof of ˜1.
+Let $\hat{\bm{Z}}=[\bm{Z}_{1},\dots,\bm{Z}_{k}]$ denote the optimal solution of Problem (4). According to Proposition˜1, it suffices to study the following two cases.
 (i) Suppose that $m<d$.
 Using this and (61), we have $\hat{\bm{Z}}_{k}=\bm{U}_{k}\bm{V}_{k}^{T}$ and $\hat{r}_{k}=m/K$ for each $k\in[K]$.
-Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (10), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=1$ and $\sum_{k=1}^{K}r_{k}=m$.
+Moreover, according to Theorem˜1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (10), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=1$ and $\sum_{k=1}^{K}r_{k}=m$.
 Since $r_{k}\leq m_{k}$ for each $k\in[K]$, we have $r_{k}=m/K$ for each $k\in[K]$.
@@
 Using this and (61), we have $\hat{\bm{Z}}_{k}=m\bm{U}_{k}\bm{V}_{k}^{T}/d$ and $\hat{r}_{k}=d/K$ for each $k\in[K]$.
-Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (9), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=m/d$ and $\sum_{k=1}^{K}r_{k}=d$.
+Moreover, according to Theorem˜1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (9), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=m/d$ and $\sum_{k=1}^{K}r_{k}=d$.
 Therefore, the global solution set of Problem (4) is a subset of that of Problem (5).
@@
 Suppose that $m,K$ are integers such that $m/K$ is a positive integer, $r\leq m/K$ is an integer, and $\alpha>0$ is a constant. Consider the following optimization problem
-$\min_{\bm{x}\in\mathbb{R}^{r}}\sum_{i=1}^{r}-\log\left(1+\alpha x_{i}\right)+\sum_{i=1}^{r}\frac{1}{K}\log(1+K\alpha x_{i})\quad\mathrm{s.t.}\ \sum_{i=1}^{r}x_{i}=\frac{m}{K},\ x_{i}\geq 0,\ \forall i\in[r].$ (66)
+$\min_{\bm{x}\in\mathbb{R}^{r}}\sum_{i=1}^{r}-\log\left(1+\alpha x_{i}\right)+\sum_{i=1}^{r}\frac{1}{K}\log(1+K\alpha x_{i})\qquad\mathrm{s.t.}\ \sum_{i=1}^{r}x_{i}=\frac{m}{K},\ x_{i}\geq 0,\ \forall i\in[r].$ (66)
 If
@@
 Then, one can verify that for all $x\geq 0$,
-$f^{\prime}(x)=-\frac{\alpha}{1+\alpha x}+\frac{\alpha}{1+K\alpha x}<0,\quad f^{\prime\prime}(x)=\frac{\alpha^{2}}{(1+\alpha x)^{2}}-\frac{K\alpha^{2}}{(1+K\alpha x)^{2}}.$ (69)
+$f^{\prime}(x)=-\frac{\alpha}{1+\alpha x}+\frac{\alpha}{1+K\alpha x}<0,\qquad f^{\prime\prime}(x)=\frac{\alpha^{2}}{(1+\alpha x)^{2}}-\frac{K\alpha^{2}}{(1+K\alpha x)^{2}}.$ (69)
 Introducing dual variables $\lambda$ associated with the constraint $\sum_{i=1}^{r}x_{i}=m/K$ and $\mu_{i}$ associated with the constraint $x_{i}\geq 0$ for each $i\in[r]$, we write the Lagrangian as follows
@@
 Then, the solution of Problem (73) is $x_{i}\in\{\overline{x},\underline{x}\}$ for each $i\in\mathcal{S}$, where
-$\overline{x}=\frac{\eta+\sqrt{\Delta}}{2K\lambda\alpha},\quad\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}.$ (76)
+$\overline{x}=\frac{\eta+\sqrt{\Delta}}{2K\lambda\alpha},\qquad\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}.$ (76)
 Now, we discuss the KKT points that could potentially be optimal solutions.
@@
 According to (69), we have $f^{\prime\prime}(x)=0$ at $\hat{x}=1/(\alpha\sqrt{K})$.
-Then, we obtain that $f^{\prime}(x)$ is strictly decreasing in $[0,\hat{x}]$ and strictly increasing in $[\hat{x},\infty]$.
+Then, we obtain that $f^{\prime}(x)$ is strictly decreasing in $0,\hat{x}$] and strictly increasing in $\hat{x}$$\infty$].
 Then, one can further verify that $\underline{x}<\hat{x}<\overline{x}$.
@@
 This, together with (78) and $l\geq 2$, yields
-$f^{\prime\prime}(x_{r-1}^{\star})<0,\quad f^{\prime\prime}(x_{r}^{\star})<0.$ (79)
-Using the second-order necessary condition for constraint optimization problems (see, e.g., [35, Theorem 12.5]) and $x_{i}^{\star}\geq 0$ for all $i\in[r]$, we obtain
+$f^{\prime\prime}(x_{r-1}^{\star})<0,\qquad f^{\prime\prime}(x_{r}^{\star})<0.$ (79)
+Using the second-order necessary condition for constraint optimization problems (see, e.g., [nocedal1999numerical, Theorem 12.5]) and $x_{i}^{\star}\geq 0$ for all $i\in[r]$, we obtain
 $\sum_{i=1}^{r}f^{\prime\prime}(x_{i}^{\star})v_{i}^{2}\geq 0,\ \forall\bm{v}\in\mathbb{R}^{r}\ \text{s.t.}\ \sum_{i=1}^{r}v_{i}=0.$ (80)
@@
 $F_{2}=(s-1)f\left(\overline{x}\right)+f\left(\underline{x}\right)\geq(s-1)f\left(\frac{m}{(s-1)K}\right)+f(\underline{x})=\frac{m}{K}h\left(\frac{m}{(s-1)K}\right)+f(\underline{x}),$ (83)
-where the inequality is because $\overline{x}\leq{m}/\left({(s-1)K}\right)$ and $f(x)$ is strictly dereasing in $[0,\infty)$. For all $x\in\left[{m/(sK)},{m}/\left({(s-1)K}\right)\right]$, we compute
+where the inequality is because $\overline{x}\leq{m}/\left({(s-1)K}\right)$ and $f(x)$ is strictly dereasing in $0,\infty)$. For all $x\in\left[{m/(sK)},{m}/\left({(s-1)K}\right)\right]$, we compute
 $h^{\prime}(x)$ $=\frac{f^{\prime}(x)x-f(x)}{x^{2}}=-\frac{1}{x}\left(\frac{\alpha}{1+\alpha x}-\frac{\alpha}{1+K\alpha x}\right)+\frac{1}{x^{2}}\left(\log(1+\alpha x)-\frac{1}{K}\log(1+K\alpha x)\right)$
@@
 According to (67), we have
-$0\displaystyle 1+\frac{\log K}{K-1}-\log(1+\alpha)=\log\left(\frac{{K}^{\frac{1}{K-1}}\exp(1)}{1+\alpha}\right)<$ 0 . (85)
+$1+\frac{\log K}{K-1}-\log(1+\alpha)=\log\left(\frac{{K}^{\frac{1}{K-1}}\exp(1)}{1+\alpha}\right)<0.$ (85)
 Using the mean-value theorem, there exists $x\in\left({m/(sK)},{m}/\left({(s-1)K}\right)\right)$ such that
@@
 $s\left(-\log\left(1+\frac{\alpha m}{sK}\right)+\frac{1}{K}\log\left(1+\frac{\alpha m}{s}\right)\right).$
-This, together with Lemma 7, implies that when the optimal solution takes the form of (77) with $s=r$, Problem (66) achieves its global minimum. Then, we complete the proof. ∎
+This, together with Lemma˜7, implies that when the optimal solution takes the form of (77) with $s=r$, Problem (66) achieves its global minimum. Then, we complete the proof. ∎
 Lemma 7.
-Consider the setting in Lemma 6 and the following function
+Consider the setting in Lemma˜6 and the following function
 $h(s):=s\left(\frac{1}{K}\log\left(1+\frac{m\alpha}{s}\right)-\log\left(1+\frac{m\alpha}{sK}\right)\right),$ (88)
@@
 ∎
-D.3 Proof of and Theorem 2
-Proof of Theorem 2.
-According to (i) of Proposition 3, if $\bm{Z}$ is a critical point but not a local maximizer, we have $\bm{Z}\in\mathcal{Z}_{2}\cup\mathcal{Z}^{c}$.
-This, together with (ii) of Proposition 3 and Proposition 4, yields that $\bm{Z}$ is a strict saddle point.
+D.3 Proof of and Theorem˜2
+Proof of Theorem˜2.
+According to (i) of ˜3, if $\bm{Z}$ is a critical point but not a local maximizer, we have $\bm{Z}\in\mathcal{Z}_{2}\cup\mathcal{Z}^{c}$.
+This, together with (ii) of ˜3 and ˜4, yields that $\bm{Z}$ is a strict saddle point.
 ∎
 Appendix E Additional Experimental Setups and Results
-In this section, we provide additional implementation details and experimental results under different parameter settings for Sections 5.1 and 5.2 in Sections E.1 and E.2, respectively.
-E.1 Implementation Details and Additional Results in Section 5.1
+In this section, we provide additional implementation details and experimental results under different parameter settings for Sections˜5.1 and 5.2 in Sections˜E.1 and E.2, respectively.
+E.1 Implementation Details and Additional Results in Section˜5.1
 Training setups.
@@
 Figure 5: Heatmap of cosine similarity between pairwise features under different settings.
-In addition to the results presented in Section 5.1, we perform additional experiments under different settings as follows.
+In addition to the results presented in Section˜5.1, we perform additional experiments under different settings as follows.
 To support our theorems, we visualize the heatmap of learned features and plot the number and magnitude of singular values in each class.
@@
 In this experiment, we consider that the number of samples in each class is same and set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=m_{2}=m_{3}=m_{4}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$.
-We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(a) and the number and magnitude of singular values in each class in Figure 6.
+We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(a) and the number and magnitude of singular values in each class in Figure˜6.
   • Experiment 2 on data with more classes.
 In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=8$, the number of samples in each class is $m_{1}=50,m_{2}=40,m_{3}=30,m_{4}=20,m_{5}=30,m_{6}=40,m_{7}=40,m_{8}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$.
-We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(b) and the number and magnitude of singular values in each class in Figure 7.
+We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(b) and the number and magnitude of singular values in each class in Figure˜7.
   • Experiment 3 on data where the dimension $d$ is larger than the number of samples $m$.
@@
 Note that in this experiment, we set the learning rate as $1$.
-We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(c) and the number and magnitude of singular values in each class in Figure 8.
+We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(c) and the number and magnitude of singular values in each class in Figure˜8.
 According to the results in Figures 5(a), 5(b), 6, and 7 of Experiments 1 and 2, we observe that when the number of samples is larger than its dimension, i.e., $m\geq d$, the learned features via the MCR2 principle are within-class compressible and between-class discriminative in both balanced and unbalanced data sets.
 Moreover, the dimension of the space spanned by these features is maximized such that $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$.
-This directly supports Theorem 1.
-By comparing the function value returned by GD and that computed by the closed-form in Theorem 1, we found that GD with random initialization will always converge to a global maximizer of Problem (5) when the data is balanced, while it will always converge to a local maximizer of Problem (5) when data is unbalanced.
-This directly supports Theorem 2.
+This directly supports Theorem˜1.
+By comparing the function value returned by GD and that computed by the closed-form in Theorem˜1, we found that GD with random initialization will always converge to a global maximizer of Problem (5) when the data is balanced, while it will always converge to a local maximizer of Problem (5) when data is unbalanced.
+This directly supports Theorem˜2.
 According to the results in Figures 5(c) and 8 of Experiment 3, we observe that when the number of samples is smaller than its dimension, i.e., $m\leq d$, the learned features via the MCR2 principle are orthogonal to each other and the dimension of each subspace is equal to the number of samples, i.e., $r_{k}=m_{k}$ for each $k\in[K]$.
-This exactly supports Theorem 1.
-Indeed, when $d\geq m$ and $r_{k}=m_{k}$ for each $k\in[K]$, it follows from Theorem 1 that $\bm{V}_{k}=\bm{I}$ and thus $\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}$ for each $k\in[K]$ for each local maximizer.
-Therefore, this also supports Theorem 2 as GD with random initialization converges to a local maximizer.
+This exactly supports Theorem˜1.
+Indeed, when $d\geq m$ and $r_{k}=m_{k}$ for each $k\in[K]$, it follows from Theorem˜1 that $\bm{V}_{k}=\bm{I}$ and thus $\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}$ for each $k\in[K]$ for each local maximizer.
+Therefore, this also supports Theorem˜2 as GD with random initialization converges to a local maximizer.
 Figure 6: The number and magnitude of singular values in each subspace in Experiment 1.
@@
 MNIST: $m=1000,K=4$
+MNIST: $m=1500,K=6$
 MNIST: $m=2000,K=8$
+MNIST: $m=2500,K=10$
 CIFAR: $m=1000,K=4$
+CIFAR: $m=1500,K=6$
 CIFAR: $m=2000,K=8$
+CIFAR: $m=2500,K=10$
 Figure 9: Heatmap of cosine similarity among learned features by training deep networks on MNIST and CIFAR-10.
@@
 In the figure, the darker pixels represent higher cosine similarity between features.
-In particular, when the $(i,j)$-th pixel is close to 00 (very light blue), the features $i$ and $j$ are approximately orthogonal.
-E.2 Implementation Details and Additional Results in Section 5.2
+In particular, when the $(i,j)$-th pixel is close to 0 (very light blue), the features $i$ and $j$ are approximately orthogonal.
+E.2 Implementation Details and Additional Results in Section˜5.2
 Network architecture and training setups for MNIST.
@@
 We train the network parameters via Adam by optimizing the MCR2 function.
-For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28].
+For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [loshchilov2016sgdr].
 We terminate the algorithm when it reaches 3000 epochs.
 More experimental results on MNIST.
-Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9.
+Besides the numerical results in Table˜1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure˜9.
 We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated.
-This supports our results in Theorem 1.
+This supports our results in Theorem˜1.
 Network architecture and training setups for CIFAR-10.
@@
 We train the network parameters via Adam by optimizing the MCR2 function.
-For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28].
+For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [loshchilov2016sgdr].
 We terminate the algorithm when it reaches 4000 epochs.
 More experimental results on CIFAR-10.
-Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9.
+Besides the numerical results in Table˜1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure˜9.
 We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated.
-This supports our results in Theorem 1.
+This supports our results in Theorem˜1.